### v1

# Neural Collapse-Inspired Multi-Label Federated Learning under Label-Distribution Skew

Abstract
Federated Learning (FL) enables collaborative model training across distributed clients while preserving data privacy. However, the performance of deep learning often deteriorates in FL due to decentralized and heterogeneous data. This challenge is further amplified in multi-label scenarios, where data exhibit complex characteristics such as label co-occurrence, inter-label dependency, and discrepancies between local and global label relationships. While most existing FL research primarily focuses on single-label classification, many real-world applications, particularly in domains such as medical imaging, often involve multi-label settings. In this paper, we address this important yet underexplored scenario in FL, where clients hold multi-label data with skewed label distributions. Neural Collapse (NC) describes a geometric structure in the latent feature space where features of each class collapse to their class mean with vanishing intra-class variance, and the class means form a maximally separated configuration. Motivated by this theory, we propose a method to align feature distributions across clients and to learn high-quality, well-clustered representations. To make the NC-structure applicable to multi-label settings, where image-level features may contain multiple semantic concepts, we introduce a feature disentanglement module that extracts semantically specific features. The clustering of these disentangled class-wise features is guided by a predefined shared NC structure, which mitigates potential conflicts between client models due to diverse local data distributions. In addition, we design regularisation losses to encourage compact clustering in the latent feature space. Experiments conducted on four benchmark datasets across eight diverse settings demonstrate that our approach outperforms existing methods, validating its effectiveness in this challenging FL scenario.
Introduction
Figure 1: Within an FL system, clients may present distinct data distributions and label relationships. For example, Client 1 primarily serves patients with sun exposure-related skin lesions, whereas Client 2 mainly handles cases involving melanoma and vascular lesions. The proposed method aims to address multi-label, label-skewed FL scenarios by disentangling class-specific features and promoting structured feature clustering in the latent feature space.
Federated Learning (FL) has emerged as a promising framework for analysing sensitive data, particularly in domains like healthcare imaging, where privacy and legal constraints prevent direct data sharing. However, most existing FL approaches are limited to basic multi-class classification tasks, failing to account for more realistic scenarios where multiple labels frequently co-occur within a single sample. For instance, many diseases are interrelated, and patients often present with multiple conditions or symptoms simultaneously, rather than a single isolated issue. This multi-label nature of real-world data is often overlooked. Another core challenge in FL is data heterogeneity across clients, as each client (e.g., hospital) may possess a uniquely skewed data distribution. This variation arises from differences in geography, population demographics, medical facilities, and clinical expertise. As a result, clients may exhibit label imbalance, including majority, minority, or even missing (vacant) classes. Under multi-label scenarios, such heterogeneity becomes even more pronounced, making collaboration across clients significantly more difficult. Figure 1 illustrates this scenario, demonstrating that hospitals may hold datasets containing various sets of dermatological conditions.
In this paper, we explore the problem of multi-label FL under label-skewed conditions, where each client’s local data is both imbalanced and divergent from the global distribution. Without directly sharing local data, the goal is to collaboratively train a global model that performs well across all target classes, despite the presence of class imbalance and missing labels on certain clients. To tackle the unique challenges posed by multi-label, label-skewed FL, we propose a novel method called Federated Neural Collapse Alignment for Multi-Label Learning (FedNCAlign-ML). Our approach introduces a representation learning framework for multi-label FL, inspired by Neural Collapse (NC) theory (Papyan et al.,, 2020), to enhance feature alignment across clients with heterogeneous label distributions. Specifically, we first encode each input image using a standard feature extractor, producing an image-level representation potentially containing multiple semantic concepts. We then leverage attention mechanisms to extract class-specific representations from this image-level representation, generating disentangled, class-wise embeddings. This effectively decomposes the complex multi-label problem into multiple single-label problems. Next, motivated by the NC phenomenon, we incorporate a predefined Equiangular Tight Frame (ETF) structure to promote consistent and discriminative feature alignment across clients, even under non-IID data distributions. The ETF structure facilitates class-independent clustering, encouraging local models to align with the global class geometry rather than overfitting to their local data distributions. Additionally, we introduce two complementary regularisation terms to enhance feature clustering in the latent feature space. A rejection loss filters out noisy signals from negative features, and a contrastive loss encourages compact intra-class clustering and clear inter-class separation. The key contributions of this paper are as follows:
  • We explore the important yet underexplored problem of multi-label FL under label-skewed conditions. To this end, we propose FedNCAlign-ML, a novel method inspired by Neural Collapse theory.
  • FedNCAlign-ML effectively adapts established representation learning techniques from the multi-class to the multi-label setting, and tackles the challenges introduced by heterogeneous local data distributions, label co-occurrence, and inter-label dependency.
  • To handle the presence of multiple semantic concepts within multi-label samples, we introduce an attention-based feature disentanglement block that extracts class-specific representations from image-level features. Additionally, we incorporate two regularisation terms, one to suppress noisy negative features and another to promote compact class-wise clustering.
  • We evaluate the proposed FedNCAlign-ML method on four benchmark datasets under eight diverse settings, comparing it to state-of-the-art FL approaches. The results show that FedNCAlign-ML consistently outperforms existing methods, achieving improvements of up to 2.87% in class-wise AUC and 7.29% in class-wise F1 score.
Related Work
Heterogeneous Federated Learning
The foundational FL algorithm, FedAvg (McMahan et al.,, 2017), performs iterative weighted averaging of locally trained model parameters. This process involves local training on client devices, aggregation at a central server, and broadcasting the updated global model back to clients. While FedAvg performs well under independently and identically distributed (IID) data, real-world scenarios often involve non-IID data distributions. These distributions are commonly observed as quantity skew, label distribution skew, and feature distribution skew, all of which present significant challenges for FL. In the context of multi-class classification, numerous methods have been proposed to address such data heterogeneity within FL settings. For instance, FedProx (Li et al.,, 2020) adds an L2 regularisation term to constrain local updates and encourage proximity to the global model. FedCurv (Shoham et al.,, 2019) uses Elastic Weight Consolidation (Kirkpatrick et al.,, 2017) to preserve important local parameters to reduce divergence across clients. SCAFFOLD (Karimireddy et al.,, 2020) introduces control variates to correct client drift. SphereFed (Dong et al.,, 2022) projects features onto a hypersphere using a fixed classifier to mitigate latent feature distribution discrepancies. FedVLS (Guo et al.,, 2025) applies global model distillation to alleviate forgetting of missing classes and regularisation to amplify minority class losses, thereby improving performance on underrepresented classes. Although these approaches have demonstrated effectiveness in the multi-class setting, their performance tends to deteriorate when directly applied to the more complex multi-label scenario, where each sample may be associated with multiple co-occurring labels. In this work, we focus on addressing the challenges of multi-label FL under quantity-skewed and label-skewed data distribution.
Multi-Label Federated Learning
Although multi-label classification is a fundamental task, it remains less explored than multi-class classification in image processing. In the context of FL, the multi-label setting introduces additional and distinct challenges compared to the multi-class scenario, including label co-occurrence, inter-label dependency, and divergence between local and global label relationships. FedLGT (Liu et al.,, 2024) addresses these challenges by focusing on modelling label correlations across clients. Building upon C-Tran (Lanchantin et al.,, 2021), it leverages frozen CLIP (Radford et al.,, 2021) text embeddings to maintain consistent label relationships and reduce divergence in label dependencies among clients. In contrast, our approach focuses on understanding the structure of the latent feature space and explicitly controlling and optimising it to enhance robustness under challenging label-skewed FL conditions.
Figure 2: Framework for the proposed FedNCAlign‑ML method for addressing multi‑label, label‑skewed FL. First, an attention‑based Label‑Aware Disentanglement Module (LADM) extracts label‑specific features from image‑level representations. Then, a predefined ETF matrix serves as both the shared classifier and the source of class-wise query embeddings, enabling consistent local training across clients. Additionally, two regularisation terms are introduced to suppress noisy negative features and encourage compact intra-class clustering in the latent feature space.
Neural Collapse-Inspired Methods
Recent research has shown that during the terminal phase of training (TPT), a phenomenon known as Neural Collapse (NC) emerges (Papyan et al.,, 2020). TPT refers to the stage when overparameterized deep networks are trained on balanced multi-class classification datasets, and the model’s performance plateaus while the training loss remains non-zero and continues to decrease. NC describes a distinctive geometric structure in the latent feature space: final-layer features collapse to their corresponding within-class means, and these class-wise prototypes align to form the vertices of a simplex Equiangular Tight Frame (ETF). As this simple yet elegant mathematical structure provides an optimal linear separable state for classification, NC-inspired classifiers have been explored in various domains, including federated learning (Li et al., 2023b, ), incremental learning (Yang et al., 2023b, ; Wei et al.,, 2025), few-shot learning (Yang et al., 2023b, ), and long-tail recognition (Gao et al.,, 2024). Despite growing interest, most existing work has focused on multi-class classification, while the equally important challenge of multi-label classification remains underexplored. To address this gap, Li et al., 2023a investigates the presence of NC in multi-label settings. They find that network training on a balanced multi-label dataset also exhibits NC but in a generalised form. In particular, features corresponding to single-label instances still conform to the Simplex ETF structure, while features associated with multi-label instances tend to approximate scaled averages of the single-label prototypes. MLC-NC (Tao et al.,, 2025) uses NC to process the long-tailed multi-label classification task in a centralised setting. However, it does not consider the missing class condition, which frequently arises in label-skewed FL scenarios across individual clients. To the best of our knowledge, our paper is the first to explore multi-label FL from an NC perspective. It opens the possibility to solve the heterogeneous multi-label FL challenge by fundamentally avoiding interference due to various local data distributions, label co-occurrence, and inter-label dependency.
Preliminaries
Problem Formulation
In a multi-label, label-skewed FL setting, a total of $K$ clients collaboratively train a shared global model. Each client $k$ has a local dataset $D_{k}=\{(x^{k}_{i},y^{k}_{i})\}_{i=1}^{N_{k}}$, containing $N_{k}$ samples. Each sample $(x^{k}_{i},y^{k}_{i})$ consists of an input image $x^{k}_{i}\in\mathbb{R}^{W\times H\times 3}$ and a corresponding multi-hot label vector $y^{k}_{i}\in\mathbb{R}^{C}$, indicating the presence of one or more of the $C$ possible classes in the image. Due to the label-skewed nature of the scenarios, each client may only have access to a subset of the total class set, denoted as $C_{k}\subseteq C$. Moreover, even for the present classes, the data distribution may differ significantly across clients. The objective of the federated system is to learn a global model that can accurately recognise all $C$ classes, without requiring clients to share their local data.
Neural Collapse
In this section, we first introduce the structure of the simplex ETF, followed by a discussion of the key properties of Neural Collapse.
Simplex Equiangular Tight Frame (ETF). A simplex ETF matrix $\mathbf{M}=\{\mathbf{m}_{c}\in\mathbb{R}^{d}\mid c=1,2,\dots,C\}$ composed of $C$ column vectors, each corresponding to a class prototype in $\mathbb{R}^{d}$. The matrix is defined as:
$\mathbf{M}=\sqrt{\frac{C}{C-1}}\mathbf{U}(\mathbf{I}_{C}-\frac{1}{C-1}\mathbf{J}_{C})$ (1)
$\mathbf{m}_{a}^{\top}\mathbf{m}_{b}=\begin{cases}-\frac{1}{C-1},&\text{if }a\neq b\\ 1,&\text{otherwise}\end{cases}$ (2)
where $\mathbf{U}\in\mathbb{R}^{d\times C}$ denotes a rotation orthogonal matrix $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{C}$. $\mathbf{I}_{C}\in\mathbb{R}^{C\times C}$ is the identity matrix. $\mathbf{J}_{C}\in\mathbb{R}^{C\times C}$ is an all-ones matrix. All vectors in the Simplex ETF have equal $\ell_{2}$ norm and identical pair-wise angles.
Variability Collapse ($\mathcal{NC}_{1}$). The intra-class variability of final-layer features collapses to zero, meaning that all feature vectors corresponding to the same class converge to their corresponding class mean.
Convergence to Simplex ETF ($\mathcal{NC}_{2}$). The class means collapse to the vertices of a simplex ETF, forming a maximally symmetric and equidistant configuration.
Self-Duality ($\mathcal{NC}_{3}$). Upon appropriate rescaling, the weights of the final classification layer also collapse to the class means. This results in a self-dual structure, where both the classifier weights and class means share the same simplex ETF structure.
Proposed Method
The discovery and analysis of NC in prior work reveal a promising structure for optimal feature clustering in the latent feature space under ideal training conditions. This insight provides valuable guidance on how to shape the latent feature space when data distributions are imperfect. In this work, we aim to leverage the desirable properties of NC within the challenging multi-label FL setting. An overview of the proposed framework is illustrated in Figure 2.
Label-Aware Disentanglement Module
Multi-label classification involves processing images that may contain multiple objects from different categories. This poses challenges for conventional image classification pipelines, which typically employ global average pooling to produce image-level feature representations. However, such pooled representations are often suboptimal for multi-label tasks because they compress multiple semantic concepts into a single vector, making it difficult to learn distinct, class-specific features in the latent feature space. Additionally, the varying number of labels per image further complicates representation learning and limits a model’s ability to structure and control the learned features effectively.
To address these challenges, we draw inspiration from recent advances in multi-label learning (Liu et al.,, 2021; Ridnik et al.,, 2023; Xu et al.,, 2023), which focus on extracting label-specific features. Building on this idea, we incorporate an attention mechanism to extract class-specific information from the image-level representation, thereby enhancing the discriminative power of each label. In our proposed pipeline, each input image is first processed by a standard feature extractor such as a CNN or transformer to produce a global feature representation. This is then passed to the Label-Aware Disentanglement Module (LADM), which utilises a 4-head cross-attention mechanism to derive label-specific features. In the FL setting, each client trains exclusively on its local data, which often deviates significantly from the global distribution. To address this issue, we adopt a predefined ETF matrix, used both as the classifier and as the source of class-wise query embeddings in LADM. By sharing this fixed query matrix across all clients, LADM ensures a consistent strategy for extracting disentangled, class-specific features during local training. This consistency reduces conflicts during model aggregation and promotes better feature alignment in the global model. The following equations detail the overall process:
$\mathbf{H}$ $=\text{LADM}(\mathbf{M},\mathbf{f})$ (3)
$=\text{Concat}_{0}\left(\text{MultiHeadAttn}(\mathbf{m}_{c},\mathbf{f})\right)_{c=1}^{C},$
where $\mathbf{f}\in\mathbb{R}^{d}$ denotes the image-level feature, serving as both the key and value in the attention mechanism. The output $\mathbf{H}\in\mathbb{R}^{C\times d}$ represents the class-specific features generated by LADM. $\mathbf{M}\in\mathbb{R}^{d\times C}$ is a predefined ETF matrix that provides class-wise query embeddings, where $\mathbf{m}_{c}$ is the query for class $c$. $\text{Concat}_{0}$ denotes concatenation along dimension 0.
Neural Collapse-Inspired Feature Alignment
After disentangling image-level features into class-specific representations, each feature captures a distinct semantic concept. This enables the effective adaptation of established representation learning techniques from the multi-class setting to the multi-label context. Rather than allowing each local model to independently learn its own classification boundaries, which can be hindered by label imbalance or missing classes, we fix the classifier using a predefined simplex ETF matrix. This design enforces maximal angular separation between class embeddings, enhancing the model’s ability to distinguish between classes regardless of the local label distribution. By reducing the influence of local label co-occurrence, this approach also helps mitigate divergence between local and global data distributions. Once class-specific features are extracted via LADM, each feature is projected onto its corresponding ETF classifier weight to compute a class-specific logit. These logits are then passed through a sigmoid activation to obtain binary predictions. The model is trained using binary cross-entropy (BCE) loss. The process is formalised as follows:
$\hat{y}_{i}^{c}=\text{Sigmoid}(\mathbf{h}^{c}_{i}\cdot\mathbf{m}_{c})$ (4)
$\mathcal{L}_{\text{BCE}}=-\frac{1}{C}\frac{1}{N}\sum_{c=1}^{C}\sum_{i=1}^{N}\left[y_{i}^{c}\log(\hat{y}_{i}^{c})+(1-y_{i}^{c})\log(1-\hat{y}_{i}^{c})\right]$ (5)
where $\mathbf{h}_{i}^{c}\in\mathbb{R}^{d}$ denotes the class-specific feature for class $c$ extracted by LADM for the ii‑th sample, and $\mathbf{m}_{c}\in\mathbb{R}^{d}$ represents the fixed ETF classifier weight corresponding to class $c$. $C$ is the total number of classes, and $N$ denotes the number of training samples available on the client.
Reducing Noise and Improving Clustering
After extracting class-wise features using LADM, each sample is associated with multiple features, corresponding to the total number of classes. Based on the ground-truth labels, a subset of these features is considered as positive, while the remaining ones are treated as negative. To promote more effective clustering in the latent feature space, we introduce two additional regularization terms. These terms are designed to suppress potential noise introduced by negative features and to encourage compact clustering of positive features belonging to the same class.
Negative Feature Rejection Loss: The $\mathcal{L}_{\text{BCE}}$ loss encourages each class-wise feature to align with its corresponding ETF prototype if it corresponds to a ground-truth (positive) label, and to be pushed away from its prototype if it corresponds to a negative label. To further reduce prediction ambiguity, it is important to prevent negative features from aligning with any other class prototypes. To address this, we compute the similarity between negative features and all class-wise prototypes, applying a penalty to high similarity scores. This regularization helps suppress the influence of negative features on the final prediction and mitigates potential noise. The process is formalised as follows:
$\hat{y}_{ij}^{-}=\text{Sigmoid}(\mathbf{h}^{-}_{i}\cdot\mathbf{m}_{j})$ (6)
$\mathcal{L}_{\text{Neg}}=-\frac{1}{C}\frac{1}{N}\sum_{j=1}^{C}\sum_{i=1}^{N}\mathbb{I}(\hat{y}_{ij}^{-}>\tau)\cdot\log(1-\hat{y}_{ij}^{-})$ (7)
where $h_{i}^{-}\in\mathbb{R}^{d}$ denotes the negative feature of sample $i$. The threshold $\tau$ is a predefined value used to filter out negative features with low similarity scores, ensuring that only confident negative predictions contribute to the loss. In our experiments, we empirically set $\tau=0.3$.
Positive Feature Contrastive Loss: To encourage more compact and discriminative class-wise clustering in the latent feature space, we introduce a contrastive loss. This loss enforces positive features to remain close to their corresponding class prototypes while being pushed away from all other class prototypes. The formulation is:
$\mathcal{L}_{\text{Pos}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{\mathbf{h}^{c}_{i}\cdot\mathbf{m}_{c}}}{\sum_{j=1}^{C}e^{\mathbf{h}^{c}_{i}\cdot\mathbf{m}_{j}}}$ (8)
where $\mathbf{h}^{c}_{i}\in\mathbb{R}^{d}$ denotes the positive feature of sample $i$ for class $c$. $\mathbf{m}_{j}$ represents the predefined prototype for class $j$.
With the two regularization terms, one designed to suppress noise from negative features and the other to encourage compact clustering of positive features, the total loss can be expressed as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{BCE}}+\lambda_{1}\mathcal{L}_{\text{Neg}}+\lambda_{2}\mathcal{L}_{\text{Pos}}$ (9)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters used to balance the BCE loss and the two regularization terms.
$\beta=0.5,\gamma=0.5$ $\beta=0.1,\gamma=0.5$
macro-AUC macro-F1 AUC F1 macro-AUC macro-F1 AUC F1
Centralized (Upperbound) 91.17 61.40 91.04 62.22 91.17 61.40 91.04 62.22
FedAvg (McMahan et al.,, 2017) 82.46 40.65 81.83 41.24 78.84 31.15 77.00 34.98
FedCurv (Shoham et al.,, 2019) 82.19 39.36 81.81 39.61 79.03 32.31 77.97 36.13
FedProx (Li et al.,, 2020) 82.23 39.37 81.68 39.27 79.40 31.01 76.69 35.73
SCAFFOLD (Karimireddy et al.,, 2020) 82.01 38.41 81.49 38.83 79.29 32.38 77.74 35.90
SphereFed (Dong et al.,, 2022) 84.85 47.58 85.04 47.29 83.27 39.52 79.45 40.77
FedLGT (Liu et al.,, 2024) 84.24 45.45 84.16 45.99 81.44 35.95 81.49 38.78
[…]
Table 1: Comparisons on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009). Non-IID data distributions across clients are simulated using the Dirichlet concentration parameter ($\beta$) and the class-missing ratio ($\gamma$).
$\beta=0.5,\gamma=0.71$ $\beta=0.1,\gamma=0.71$
macro-AUC macro-F1 AUC F1 macro-AUC macro-F1 AUC F1
Centralized (Upperbound) 91.86 63.49 94.52 73.30 91.86 63.49 94.52 73.30
FedAvg (McMahan et al.,, 2017) 90.18 53.96 92.68 67.53 83.95 40.69 86.68 63.03
FedCurv (Shoham et al.,, 2019) 90.31 56.17 92.64 68.11 86.20 42.19 89.54 64.76
FedProx (Li et al.,, 2020) 90.16 55.85 92.44 67.95 83.73 42.98 87.96 63.50
SCAFFOLD (Karimireddy et al.,, 2020) 89.56 54.88 92.13 67.59 85.33 43.60 89.41 64.75
SphereFed (Dong et al.,, 2022) 85.36 44.25 89.46 53.40 83.37 41.90 88.92 50.94
FedLGT (Liu et al.,, 2024) 87.33 56.24 90.84 68.25 84.80 44.09 88.70 57.35
[…]
Table 2: Comparisons on the multi-label DermaMNIST dataset (Yang et al., 2023a, ).
$\beta=0.05,\gamma=0.5$ $\beta=0.01,\gamma=0.5$
macro-AUC macro-F1 AUC F1 macro-AUC macro-F1 AUC F1
Centralized (Upperbound) 95.66 74.77 96.22 77.15 95.66 74.77 96.22 77.15
FedAvg (McMahan et al.,, 2017) 93.72 57.00 94.20 64.62 93.29 49.16 93.90 61.08
FedCurv (Shoham et al.,, 2019) 93.69 57.32 94.40 64.74 93.33 48.93 93.87 61.35
FedProx (Li et al.,, 2020) 93.70 57.08 94.26 64.46 93.05 48.99 93.74 61.16
SCAFFOLD (Karimireddy et al.,, 2020) 93.07 58.20 93.97 65.86 93.44 50.89 93.93 62.10
SphereFed (Dong et al.,, 2022) 85.72 33.71 87.38 38.19 86.85 32.44 87.94 36.18
FedLGT (Liu et al.,, 2024) 91.07 61.40 91.39 65.88 91.20 56.59 91.20 64.58
[…]
Table 3: Comparison of related FL methods with ours on the multi-label PASCAL VOC 2012 dataset (Everingham et al.,, 2010).
Experiment
Dataset and Evaluation Metric
Datasets: We conduct experiments on both general computer vision (CV) datasets and medical imaging datasets to evaluate the effectiveness and real-world applicability of the proposed method. For general CV tasks, we use CIFAR-10 (Krizhevsky et al.,, 2009) and PASCAL VOC (Everingham et al.,, 2010). Since CIFAR-10 is originally a multi-class dataset, we adopt the method used by Li et al., 2023a to transform it into a multi-label dataset. This is done by merging multiple images into a single composite image, with the combined labels serving as the new ground truth. For medical imaging tasks, we use DermaMNIST (Yang et al., 2023a, ) and ChestX-ray14 (Wang et al.,, 2017). DermaMNIST, which originally contains 7 skin disease categories in a multi-class format, is converted to a multi-label dataset using the same strategy applied to CIFAR-10. ChestX-ray14 is naturally a multi-label dataset, comprising 14 thoracic disease categories and an additional “No Finding” label. Since a significant portion of the dataset, approximately four-sevenths of the training data, are “No Finding” samples (negative cases with all-zero labels), we distribute these samples evenly across all clients. This setup mimics a realistic clinical scenario in which healthy cases are prevalent, while disease cases are relatively rare and unevenly distributed across clients. Detailed information about the datasets and local data distributions under various experimental settings is provided in the Appendix.
Evaluation Metric: Given our focus on label-skewed data distributions, where certain classes may be missing or underrepresented, we report both overall (micro) and class-wise (macro) performance metrics. Following standard practices across different datasets, we report AUC and F1 scores for CIFAR-10, PASCAL VOC, and DermaMNIST, and AUC for ChestX-ray14. Together, these metrics offer a balanced and comprehensive evaluation by capturing both overall performance and class-specific behaviour.
Task Setup and Implementation Details
Task Setup: We simulate the FL system with 10 clients and adopt a full client participation strategy. To reflect the standard heterogeneous FL setting, we introduce non-IID data distributions by partitioning the data using a Dirichlet distribution, parameterised by the concentration factor $\beta$. To further emulate label-skewed scenarios, we define a class presence ratio $\gamma$, which limits the number of classes available to each client and simulates cases where certain classes are entirely absent from local datasets.
Implementation Details: For all experiments, we use ResNet-18 (He et al.,, 2016) as the feature extractor. Each FL global model is trained over 100 communication rounds, with one epoch of local training per round. The final model is selected based on its best performance on the validation set. We use a batch size of 32 and initialise the learning rate at $1\times 10^{-4}$, using the AdamW optimiser with a weight decay of 0.01. Models for the CIFAR-10 dataset are trained from scratch, while models for other datasets are initialised with ImageNet-pretrained weights. During training, the regularisation hyperparameter $\lambda_{1}$ is empirically set to either 1 or 0.01, depending on the background complexity of the dataset. For datasets with complex backgrounds, where disentangling class-specific features is more challenging, a smaller value ($\lambda_{1}=0.01$) is used to reduce potential interference with positive feature learning. The other regularisation hyperparameter, $\lambda_{2}$, is fixed at 1 across all experiments.
Performance Comparison
To evaluate the effectiveness of the proposed method, we compare it with other methods across four datasets, as shown in Tables 3, 3, 3, and 4. As observed in these tables, our method consistently achieves the best performance in most cases across different datasets. The improvement is particularly notable on naturally long-tailed datasets such as DermaMNIST and ChestX-ray14, where our method yields substantial gains in class-wise performance, indicating its ability to produce more balanced predictions. Specifically, under the $\beta=0.1$ setting on DermaMNIST, our method outperforms the second-best approach by 1.49% in class-wise AUC and 7.29% in class-wise F1 score. For PASCAL VOC, when each client is limited to a maximum of half the total classes, our method achieves improvements of 2.94% and 5.49% in class-wise F1 score under the $\beta=0.05$ and $\beta=0.01$ settings, respectively. For ChestX-ray14 under the $\beta=0.5$ and $\beta=0.1$ settings, our method improves class-wise AUC by 0.64% and 1.31%, respectively. Overall, across a range of non-IID configurations with limited class availability per client, our method consistently outperforms existing FL approaches, demonstrating strong robustness to various label-skewed data distributions.
Ablation Study
To evaluate the effectiveness of each component in the proposed method, we conducted an ablation study on the multi-label DermaMNIST dataset, with the client data distribution configured using $\beta=0.1$ and $\gamma=0.71$. The experimental results are presented in Table 5. As shown in the table, incorporating LADM for class-specific feature extraction, together with the predefined ETF classifier to encourage class-wise clustering alignment across clients in the latent feature space, yields improvements of 0.43% in class-wise AUC and 7.34% in class-wise F1 score. Additional performance gains are achieved by introducing regularisation terms that enhance the model’s discriminative ability, resulting in further improvements of 3.31% in class-wise AUC and 3.43% in class-wise F1 score.
(a) Learnable FC classifier
(b) LADM + fixed ETF classifier
(c) LADM + fixed ETF + Dual regularization
Figure 3: t-SNE visualisation of test data feature embeddings on the multi-label DermaMNIST experiment with $\beta=0.1$, $\gamma=0.71$. Each colour represents a class. Observing from subfigure (a), without feature disentanglement (LADM), the model relies on undesired information, such as the number of labels per sample, for clustering.
$\beta=0.5,\gamma=0.5$ $\beta=0.1,\gamma=0.5$
macro-AUC AUC macro-AUC AUC
Centralized 71.44 79.62 71.44 79.62
FedAvg 69.07 73.28 69.84 78.00
FedCurv 68.12 72.15 69.46 77.63
FedProx 69.23 72.45 69.23 77.69
SCAFFOLD 69.02 72.21 66.71 78.69
SphereFed 57.48 70.26 61.80 74.59
FedLGT 69.76 72.41 70.02 77.93
[…]
Table 4: Comparison of related FL methods with ours on the ChestX-ray14 dataset (Wang et al.,, 2017).
ETF Clf LADM $\mathcal{L}_{\text{Neg}}$ $\mathcal{L}_{\text{Pos}}$ macro-AUC macro-F1 AUC F1
83.95 40.69 86.68 63.03
84.38 47.95 86.70 60.49
84.73 45.06 89.71 62.89
85.20 49.84 89.03 61.06
87.69 51.38 90.36 63.26
Table 5: Ablation study of the proposed method on the multi-label DermaMNIST dataset with $\beta=0.1$, $\gamma=0.71$.
query type query init macro-AUC macro-F1 AUC F1
learnable random 82.94 47.94 86.33 57.37
learnable ETF 85.39 46.92 84.94 53.37
fixed ETF 84.38 47.95 86.70 60.49
Table 6: Ablation study of the class-wise feature extraction block - LADM on the multi-label DermaMNIST dataset.
(a) FC classifier
(b) LADM + fixed ETF classifier
(c) LADM + ETF + Dual regularization
Figure 4: Pair-wise cosine similarity between test data class-wise average feature prototypes.
We also investigated the class-wise query mechanism within the attention-based LADM module. As reported in Table 6, fixed, well-designed class-wise queries consistently outperform learnable queries across most evaluation metrics. This outcome is intuitive in the context of label-skewed FL, where each client has a unique local data distribution. If the queries and classifier are learnable, the model may overfit to these local distributions, which can significantly deviate from the global distribution. Such overfitting introduces inconsistencies among client models and ultimately degrades the performance of the aggregated global model. To maintain consistent learning across clients with non-iid data, it is therefore preferable to use predefined, well-separated, fixed embeddings to guide the learning process rather than relying on learnable representations.
To further analyse model behaviour, we present t-SNE visualisations of feature distributions from the test data in the latent feature space, generated under different model architectures and training strategies on the multi-label DermaMNIST dataset. The client data distribution is configured using $\beta=0.1$ and $\gamma=0.71$. As shown in Figure 3, when a conventional, learnable fully connected (FC) classifier is used, although the model achieves 83.95% class-wise AUC and 40.69% F1 score, the resulting feature representations exhibit poor clustering. Notably, the model appears to rely on undesired information. Specifically, it tends to group features based on the number of labels per sample, rather than focusing solely on semantic content. By incorporating LADM for extracting single-class features and employing a predefined ETF classifier to regulate feature distribution across clients, the model learns to focus more on semantic content for clustering. This leads to more disentangled, class-specific feature representations and improved clustering quality. Figure 4 also shows that the pairwise cosine similarity between class-wise average features is substantially reduced, indicating enhanced inter-class separability and stronger discriminative capability. Finally, with the addition of extra regularisation terms during training, the feature clustering becomes more compact and semantically coherent, and the similarity among class-wise average prototypes is further reduced.
Conclusion
This paper investigates multi-label label-skewed FL, a practically important yet underexplored problem setting. This task is particularly challenging due to the complexities inherent in multi-label data, combined with difficulties introduced by the federated system, including heterogeneous local data distributions and inconsistencies between local and global data distributions. Inspired by Neural Collapse theory, we propose a feature disentanglement module and incorporate a predefined ETF structure to facilitate the extraction and clustering of class-specific representations, while also aligning feature distributions across clients in the latent feature space. Furthermore, we incorporate specific regularization losses to suppress noisy negative features and enhance compact class-wise clustering. Experiments on four benchmark datasets across eight settings show that our method outperforms existing approaches, with gains of up to 2.87% in class-wise AUC and 7.29% in class-wise F1 score.
Acknowledgments
This work was supported by the UKRI grant EP/X040186/1 (Turing AI Fellowship). This work was also partly supported by the InnoHK-funded Hong Kong Centre for Cerebrocardiovascular Health Engineering (COCHE) Project 2.1 (Cardiovascular risks in early life and fetal echocardiography).
Appendix
Pseudo Code for FedNCAlign-ML
We present the pseudocode for our proposed FedNCAlign-ML in Algorithm 1. The input comprises a set of clients, each with its own local dataset, and a centralized server responsible for managing communication. The training process involves three iterative steps: local training on client devices, aggregation at the central server, and broadcasting the updated global model back to the clients. While we adopt the standard FL protocol for aggregation and broadcasting, the core contribution of FedNCAlign-ML lies in its specialized model design and the local training procedure tailored for multi-label label-skewed FL scenarios. During local training, each client first encodes its input data using a standard feature extractor to generate image-level representations. These representations are then passed to the proposed Label Aware Disentanglement Module (LADM), which produces disentangled, class-wise embeddings. By capturing distinct semantic concepts within each feature, these embeddings facilitate more effective and controllable clustering in the latent feature space. To further enhance consistent and discriminative feature alignment across clients, particularly under label-skewed conditions, we introduce a shared, predefined Equiangular Tight Frame (ETF) structure. This structure encourages class-independent clustering and aligns local features with a global class geometry, helping to mitigate overfitting to local data distributions. Additionally, two complementary regularization terms are introduced to enhance feature clustering in the latent feature space: a rejection loss to filter out noisy signals from negative features, and a contrastive loss to promote compact intra-class clustering and clear inter-class separation.
Algorithm 1 Federated Neural Collapse Alignment for Multi-Label Learning (FedNCAlign-ML)
Input: $K$ clients; local datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$; initial global parameters $w_{0}$; predefined ETF matrix $\mathbf{M}$; learning rate $\eta$; local epochs $E$.
1:Server:
2:Initialize $w\leftarrow w_{0}$
3:for $t=0$ to $T$ do $\triangleright$ $t$-th communication round
4:  for each client $k=1,\dots,K$ in parallel do
5:   $w_{t+1}^{k}\leftarrow\textsc{ClientUpdate}(\mathcal{D}_{k},w_{t})$
6:   Send the client model $w^{k}_{t+1}$ to the server
7:  end for
8:  $w_{t+1}\leftarrow\textsc{ModelAggregation}(\{w_{t+1}^{k}\}_{k=1}^{K})$
9:  Broadcast the global model $w_{t+1}$ to all clients
10:end for
11:ClientUpdate($\mathcal{D}_{k},w$):
12:Split $\mathcal{D}_{k}$ into batches $\{b\}$
13:for $e=0$ to $E$ do $\triangleright$ $e$-th local epoch
14:  for each batch $b$ in $\{b\}$ do
15:   Extract inputs $(x,y)$ from $b$
16:   $\mathbf{f}\leftarrow\text{FeatureExtractor}(w;x)$
17:   $\mathbf{H}\leftarrow\text{LADM}(\mathbf{M},\mathbf{f})$ $\triangleright$ Eq. 3
18:   Compute loss $\mathcal{L}_{\text{total}}(w;\mathbf{M},\mathbf{H})$ $\triangleright$ Eqs. 5789
19:   $w\leftarrow w-\eta\nabla_{w}\mathcal{L}_{\text{total}}$
20:  end for
21:end for
22:Return $w$
Ablation Study (additional)
To complement the results reported in the main manuscript, we present an additional ablation study to evaluate the contribution of each component in our proposed method. The experiments are conducted on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009), with the client data distribution configured using a Dirichlet concentration parameter $\beta=0.5$ and a class-presenting ratio $\gamma=0.5$. The experimental results are summarized in Tables 7 and 8, and Figures 5 and 6. As shown in Table 7, integrating LADM for class-specific feature extraction, along with the use of a predefined ETF classifier to encourage class-wise clustering alignment across clients in the latent feature space, leads to performance improvements of 3.38% in class-wise AUC and 3.99% in class-wise F1 score. Additional gains are achieved by introducing regularization terms that enhance the model’s discriminative capacity, resulting in a further class-wise increase of 1.88% in AUC and 4.58% in F1 score. Regarding LADM specifically, the results in Table 8 demonstrate that using fixed, well-designed class-wise queries is generally more effective than learnable queries across most evaluation metrics.
To further analyze model behavior, we visualize the latent feature distributions of the test set using t-SNE under different architectural and training configurations. As shown in Figure 5, incorporating LADM for extracting single-class features, along with a predefined ETF classifier to regulate feature distribution across clients, enables the model to focus on semantic content rather than irrelevant factors such as the number of labels per sample. Furthermore, the addition of regularization terms during training leads to more compact and semantically coherent clusters, while also reducing the similarity among class-wise average prototypes. As shown in Figure 6, the pairwise cosine similarity between class-wise average features decreases with the inclusion of LADM and the ETF classifier, and is further reduced by the added regularization terms, indicating enhanced inter-class separability and stronger discriminative capability.
ETF Clf LADM $\mathcal{L}_{\text{Neg}}$ $\mathcal{L}_{\text{Pos}}$ macro-AUC macro-F1 AUC F1
82.46 40.65 81.83 41.24
85.84 44.64 85.10 45.45
87.72 48.22 86.16 48.93
86.78 44.46 86.08 45.28
87.72 49.22 87.13 49.60
Table 7: Ablation study of the proposed method on the multi-label CIFAR10 dataset. $\beta$ is set to 0.5 and $\gamma$ is set to 0.5.
query type query init macro-AUC macro-F1 AUC F1
learnable random 84.15 38.10 83.70 39.30
learnable ETF 86.26 43.61 85.41 43.87
fixed ETF 85.84 44.64 85.10 45.45
Table 8: Ablation study of the class-wise feature extraction block - LADM on the multi-label CIFAR10 dataset.
(a) Learnable FC classifier
(b) LADM + fixed ETF classifier
(c) LADM + fixed ETF + Dual regularization
Figure 5: t-SNE visualisation of test data feature embeddings on the multi-label CIFAR10 experiment with $\beta=0.5$, $\gamma=0.5$. Each colour represents a class. Observing from subfigure (a), without feature disentanglement (LADM), the model appears to rely on undesired information, such as the number of labels per sample, for clustering.
(a) FC classifier
(b) LADM + fixed ETF classifier
(c) LADM + ETF + Dual regularization
Figure 6: Pair-wise cosine similarity between test data class-wise average feature prototypes on the multi-label CIFAR10 dataset.
$\beta=0.5,\gamma=0.5$
macro-AUC macro-F1 AUC F1
Centralized (Upperbound) $92.20\pm 0.36$ $61.30\pm 0.94$ $90.04\pm 0.65$ $62.02\pm 0.62$
FedAvg (McMahan et al.,, 2017) $82.29\pm 0.46$ $39.47\pm 1.59$ $81.48\pm 0.78$ $40.26\pm 1.09$
FedCurv (Shoham et al.,, 2019) $82.53\pm 0.30$ $39.96\pm 1.28$ $82.10\pm 0.29$ $40.34\pm 1.04$
FedProx (Li et al.,, 2020) $82.45\pm 0.34$ $39.22\pm 0.74$ $81.77\pm 0.48$ $39.66\pm 0.53$
SCAFFOLD (Karimireddy et al.,, 2020) $82.51\pm 0.44$ $39.98\pm 1.41$ $82.08\pm 0.53$ $40.26\pm 1.26$
SphereFed (Dong et al.,, 2022) $\underline{83.63}\pm 1.50$ $42.58\pm 3.20$ $\underline{83.50}\pm 1.87$ $43.18\pm 2.34$
FedLGT (Liu et al.,, 2024) $83.52\pm 0.68$ $\underline{43.60}\pm 1.68$ $83.36\pm 0.71$ $\underline{44.03}\pm 1.71$
[…]
Table 9: Comparisons on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009). Non-IID data distributions across clients are simulated using the Dirichlet concentration parameter ($\beta$) and the class-missing ratio ($\gamma$).
Experiments (additional)
To complement the results reported in the main manuscript and further validate the reproducibility of our method, we present additional experiments on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009). These experiments were repeated three times with different random seeds to evaluate the stability and effectiveness of the proposed approach. The mean and standard deviation of the results are reported in Table 9. As shown in the table, the proposed method demonstrates stable performance across multiple runs and consistently outperforms competing approaches.
Dataset
In this section, we introduce the datasets used in our experiments. To evaluate both the effectiveness and real-world applicability of the proposed method, we conduct experiments on datasets from general computer vision (CV) as well as medical imaging domains. Specifically, we use CIFAR-10 (Krizhevsky et al.,, 2009) and PASCAL VOC (Everingham et al.,, 2010) as general CV datasets, and DermaMNIST (Yang et al., 2023a, ) and ChestX-ray14 (Wang et al.,, 2017) as medical datasets. We simulate the FL system with 10 clients. To model non-IID data distributions, the data is partitioned using a Dirichlet distribution, with the concentration parameter $\beta$ controlling the degree of heterogeneity. To further simulate missing-class scenarios, we constrain the number of classes available to each client using the class-presenting ratio $\gamma$, which specifies the proportion of total classes present locally. Dataset-specific settings are detailed below.
CIFAR-10 (Krizhevsky et al.,, 2009) is a widely used benchmark in CV. It comprises 10 classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. To adapt CIFAR-10 for multi-label classification, following Li et al., 2023a , we create composite samples by combining multiple original images into a single image, with their corresponding labels forming a multi-label ground truth. Specifically, we retain a portion of the original single-label images and augment the dataset by creating an equal number of synthetic samples for each possible pairwise class combination. These composites are constructed by randomly selecting two images from different classes and combining them into one. We evaluate our method on the multi-label CIFAR-10 dataset under two FL configurations, using $\beta=0.5,\ \gamma=0.5$ and $\beta=0.1,\ \gamma=0.5$. Figure 8 illustrates the corresponding class-wise data distribution across clients.
DermaMNIST (Yang et al., 2023a, ) is a skin lesion classification dataset. It consists of 7 diagnostic categories: actinic keratoses, basal cell carcinoma, benign keratosis-like lesions, dermatofibroma, melanocytic nevi, melanoma, and vascular lesions. To adapt DermaMNIST for multi-label classification, we apply a strategy similar to the one used for multi-label CIFAR-10. Given the long-tailed nature of the original dataset, we preserve its inherent distribution by retaining all single-label samples. Then, we augment the dataset with an equal number of synthetic samples for each possible pairwise label combination. Figure 7 shows examples of the resulting composite multi-label inputs. The resulting dataset comprises the complete set of original samples along with the newly generated multi-label samples. We evaluate our method on this multi-label DermaMNIST dataset under two FL configurations: $\beta=0.5,\gamma=0.71$ and $\beta=0.1,\gamma=0.71$. The corresponding class-wise data distributions across clients are illustrated in Figure 9.
(a) actinic keratoses, vascular lesions
(b) basal cell carcinoma, dermatofibroma
(c) bird, vascular lesions
(d) dermatofibroma, melanoma
Figure 7: Examples from the multi-label DermaMNIST dataset. Each subfigure displays a composite image along with its corresponding set of labels.
PASCAL VOC 2012 (VOC) (Everingham et al.,, 2010) is a widely used benchmark dataset in CV. It consists of approximately 11,500 images, each annotated with one or more object categories selected from a predefined set of 20 classes. These object categories are: aeroplane, bicycle, bird, boat, bottle, bus, car, cat, chair, cow, dining table, dog, horse, motorbike, person, potted plant, sheep, sofa, train, and TV monitor. This dataset poses a challenging multi-label classification problem due to the presence of class imbalance, intra-class variability, and frequent inter-class co-occurrence. For example, the person category often co-occurs with many other object classes. Under a label-skewed FL setting, the non-IID distribution of data and inconsistencies in class co-occurrence patterns across clients further exacerbate the learning difficulty. We evaluate our method on VOC under two FL configurations, using $\beta=0.05$, $\gamma=0.5$ and $\beta=0.01$, $\gamma=0.5$. Figure 10 illustrates the resulting class-wise data distribution across clients.
ChestX-ray14 (Wang et al.,, 2017) is a large-scale medical imaging dataset commonly used for automated thoracic disease classification. It consists of 112,120 frontal-view chest X-ray images collected from 30,805 unique patients. Each image is annotated with one or more disease labels, making the dataset naturally suitable for multi-label classification tasks. These labels were extracted from corresponding radiology reports using natural language processing techniques. The dataset includes 14 disease categories: Atelectasis, Cardiomegaly, Consolidation, Edema, Effusion, Emphysema, Fibrosis, Hernia, Infiltration, Mass, Nodule, Pleural Thickening, Pneumonia, and Pneumothorax. In addition, a No Finding label is used to indicate negative samples in which none of the 14 diseases are present. Due to its scale, real-world variability, and label noise, ChestX-ray14 provides a challenging benchmark for developing and evaluating deep learning models in medical image analysis, particularly in multi-label, imbalanced scenarios. We conduct experiments under two label-skewed FL settings on this dataset, using configurations of $\beta=0.5$, $\gamma=0.5$ and $\beta=0.1$, $\gamma=0.5$. Since a significant portion of the dataset is “No Finding” samples, we distribute these evenly across all clients in both settings. This setup mimics realistic clinical scenarios where healthy cases are common, while disease cases are relatively rare and unequally distributed across institutions. Figure 11 shows the class-wise data distribution across clients.
(a) Overall training data distribution.
(b) Per-client training data distribution with $\beta=0.5$ and $\gamma=0.5$.
(c) Per-client training data bubble chart under $\beta=0.5$ and $\gamma=0.5$.
(d) Testing data distribution.
(e) Per-client training data distribution under $\beta=0.1$ and $\gamma=0.5$.
(f) Per-client training data bubble chart under $\beta=0.1$ and $\gamma=0.5$.
Figure 8: Distribution of data across local clients used in the experiments on the CIFAR-10 dataset (Krizhevsky et al.,, 2009).
(a) Overall training data distribution.
(b) Per-client training data distribution under $\beta=0.5$ and $\gamma=0.71$.
(c) Per-client training data bubble chart under $\beta=0.5$ and $\gamma=0.71$.
(d) Testing data distribution.
(e) Per-client training data distribution under $\beta=0.1$ and $\gamma=0.71$.
(f) Per-client training data bubble chart under $\beta=0.1$ and $\gamma=0.71$.
Figure 9: Distribution of data across local clients used in the experiments on the DermaMNIST dataset (Yang et al., 2023a, ).
(a) Overall training data distribution.
(b) Per-client training data distribution under $\beta=0.05$ and $\gamma=0.5$.
(c) Per-client training data bubble chart under $\beta=0.05$ and $\gamma=0.5$.
(d) Testing data distribution.
(e) Per-client training data distribution under $\beta=0.01$ and $\gamma=0.5$.
(f) Per-client training data bubble chart under $\beta=0.01$ and $\gamma=0.5$.
Figure 10: Distribution of data across local clients used in the experiments on the VOC dataset (Everingham et al.,, 2010).
(a) Overall training data distribution.
(b) Per-client training data distribution under $\beta=0.5$ and $\gamma=0.5$.
(c) Per-client training data bubble chart under $\beta=0.5$ and $\gamma=0.5$.
(d) Testing data distribution.
(e) Per-client training data distribution under $\beta=0.1$ and $\gamma=0.5$.
(f) Per-client training data bubble chart under $\beta=0.1$ and $\gamma=0.5$.
Figure 11: Distribution of data across local clients used in the experiments on the ChestX-ray14 dataset (Wang et al.,, 2017).

### v2

@@
 Problem Formulation
-In a multi-label, label-skewed FL setting, a total of $K$ clients collaboratively train a shared global model.
-Each client $k$ has a local dataset $D_{k}=\{(x^{k}_{i},y^{k}_{i})\}_{i=1}^{N_{k}}$, containing $N_{k}$ samples.
-Each sample $(x^{k}_{i},y^{k}_{i})$ consists of an input image $x^{k}_{i}\in\mathbb{R}^{W\times H\times 3}$ and a corresponding multi-hot label vector $y^{k}_{i}\in\mathbb{R}^{C}$, indicating the presence of one or more of the $C$ possible classes in the image.
-Due to the label-skewed nature of the scenarios, each client may only have access to a subset of the total class set, denoted as $C_{k}\subseteq C$.
-Moreover, even for the present classes, the data distribution may differ significantly across clients.
-The objective of the federated system is to learn a global model that can accurately recognise all $C$ classes, without requiring clients to share their local data.
-Neural Collapse
+We consider a multi-label, label-skewed FL setting with $K$ clients collaboratively training a shared global model.
+Each client $k$ holds a private dataset $\mathcal{D}_{k}=\{(x_{i}^{k},y_{i}^{k})\}_{i=1}^{N_{k}}$ of size $N_{k}$, where the input image $x_{i}^{k}\in\mathbb{R}^{H\times W\times 3}$ and the label $y_{i}^{k}\in\{0,1\}^{C}$ is a multi-hot vector over the $C$ classes.
+Let $\mathcal{C}=\{1,\dots,C\}$ denote the global class index set.
+Due to label skew, client $k$ may only observe a subset $\mathcal{C}_{k}\subseteq\mathcal{C}$, and even for shared classes, the class-conditional distributions can differ across clients.
+The goal is to learn a single global model that accurately recognizes all classes in $\mathcal{C}$ without sharing raw data among clients.
+Neural Collapse (NC)
 In this section, we first introduce the structure of the simplex ETF, followed by a discussion of the key properties of Neural Collapse.
 Simplex Equiangular Tight Frame (ETF).
-A simplex ETF matrix $\mathbf{M}=\{\mathbf{m}_{c}\in\mathbb{R}^{d}\mid c=1,2,\dots,C\}$ composed of $C$ column vectors, each corresponding to a class prototype in $\mathbb{R}^{d}$.
-The matrix is defined as:
-$\mathbf{M}=\sqrt{\frac{C}{C-1}}\mathbf{U}(\mathbf{I}_{C}-\frac{1}{C-1}\mathbf{J}_{C})$ (1)
-$\mathbf{m}_{a}^{\top}\mathbf{m}_{b}=\begin{cases}-\frac{1}{C-1},&\text{if }a\neq b\\ 1,&\text{otherwise}\end{cases}$ (2)
-where $\mathbf{U}\in\mathbb{R}^{d\times C}$ denotes a rotation orthogonal matrix $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{C}$.
-$\mathbf{I}_{C}\in\mathbb{R}^{C\times C}$ is the identity matrix.
-$\mathbf{J}_{C}\in\mathbb{R}^{C\times C}$ is an all-ones matrix.
-All vectors in the Simplex ETF have equal $\ell_{2}$ norm and identical pair-wise angles.
-Variability Collapse ($\mathcal{NC}_{1}$).
-The intra-class variability of final-layer features collapses to zero, meaning that all feature vectors corresponding to the same class converge to their corresponding class mean.
-Convergence to Simplex ETF ($\mathcal{NC}_{2}$). The class means collapse to the vertices of a simplex ETF, forming a maximally symmetric and equidistant configuration.
-Self-Duality ($\mathcal{NC}_{3}$).
-Upon appropriate rescaling, the weights of the final classification layer also collapse to the class means.
-This results in a self-dual structure, where both the classifier weights and class means share the same simplex ETF structure.
+A simplex ETF matrix $\mathbf{M}=[\mathbf{m}_{c}]_{c=1}^{C}\in\mathbb{R}^{d\times C}$ composed of $C$ column vectors, each corresponding to a class prototype in $\mathbf{m}_{c}\in\mathbb{R}^{d}$.
+A standard construction is:
+$\mathbf{M}\;=\;\sqrt{\frac{C}{C-1}}\;\mathbf{U}\;\Big(\mathbf{I}_{C}-\frac{1}{C}\,\mathbf{1}_{C}\mathbf{1}_{C}^{\top}\Big),$ (1)
+where $\mathbf{U}\in\mathbb{R}^{d\times C}$ denotes a rotation orthogonal matrix $(\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{C})$.
+$\mathbf{I}_{C}$ is the $C\times C$ identity matrix.
+$\mathbf{1}_{C}$ is the $C$-dimensional all-ones vector.
+This yields unit-norm columns with equal pairwise inner products:
+$\mathbf{m}_{a}^{\top}\mathbf{m}_{b}=\begin{cases}1,&a=b,\\[2.0pt] -\dfrac{1}{C-1},&a\neq b,\end{cases}\qquad a,b\in\mathcal{C}.$ (2)
+Hence, all prototypes have equal $\ell_{2}$ norm and identical pairwise angles, forming a centered regular simplex.
+Neural Collapse properties.
+At the end of training, over-parameterized networks exhibit the NC phenomenon. Empirically, the following regularities are observed:
+  • Variability Collapse ($\mathcal{NC}_{1}$). Within-class feature variance collapses: features of the same class concentrate at their class mean.
+  • Convergence to Simplex ETF ($\mathcal{NC}_{2}$). Class means arrange as the vertices of a simplex ETF, forming a maximally symmetric and equidistant configuration.
+  • Self-Duality ($\mathcal{NC}_{3}$). Upon appropriate rescaling, the final-layer classifier weights align with the class means, sharing the same simplex ETF geometry.
 Proposed Method
-The discovery and analysis of NC in prior work reveal a promising structure for optimal feature clustering in the latent feature space under ideal training conditions.
-This insight provides valuable guidance on how to shape the latent feature space when data distributions are imperfect.
-In this work, we aim to leverage the desirable properties of NC within the challenging multi-label FL setting.
-An overview of the proposed framework is illustrated in Figure 2.
-Label-Aware Disentanglement Module
-Multi-label classification involves processing images that may contain multiple objects from different categories.
-This poses challenges for conventional image classification pipelines, which typically employ global average pooling to produce image-level feature representations.
-However, such pooled representations are often suboptimal for multi-label tasks because they compress multiple semantic concepts into a single vector, making it difficult to learn distinct, class-specific features in the latent feature space.
-Additionally, the varying number of labels per image further complicates representation learning and limits a model’s ability to structure and control the learned features effectively.
-To address these challenges, we draw inspiration from recent advances in multi-label learning (Liu et al.,, 2021; Ridnik et al.,, 2023; Xu et al.,, 2023), which focus on extracting label-specific features.
-Building on this idea, we incorporate an attention mechanism to extract class-specific information from the image-level representation, thereby enhancing the discriminative power of each label.
-In our proposed pipeline, each input image is first processed by a standard feature extractor such as a CNN or transformer to produce a global feature representation.
-This is then passed to the Label-Aware Disentanglement Module (LADM), which utilises a 4-head cross-attention mechanism to derive label-specific features.
-In the FL setting, each client trains exclusively on its local data, which often deviates significantly from the global distribution.
-To address this issue, we adopt a predefined ETF matrix, used both as the classifier and as the source of class-wise query embeddings in LADM.
-By sharing this fixed query matrix across all clients, LADM ensures a consistent strategy for extracting disentangled, class-specific features during local training.
-This consistency reduces conflicts during model aggregation and promotes better feature alignment in the global model.
-The following equations detail the overall process:
-$\mathbf{H}$ $=\text{LADM}(\mathbf{M},\mathbf{f})$ (3)
-$=\text{Concat}_{0}\left(\text{MultiHeadAttn}(\mathbf{m}_{c},\mathbf{f})\right)_{c=1}^{C},$
-where $\mathbf{f}\in\mathbb{R}^{d}$ denotes the image-level feature, serving as both the key and value in the attention mechanism.
-The output $\mathbf{H}\in\mathbb{R}^{C\times d}$ represents the class-specific features generated by LADM.
-$\mathbf{M}\in\mathbb{R}^{d\times C}$ is a predefined ETF matrix that provides class-wise query embeddings, where $\mathbf{m}_{c}$ is the query for class $c$.
-$\text{Concat}_{0}$ denotes concatenation along dimension 0.
+The NC phenomenon suggests a promising structure for optimal feature clustering in the latent feature space under well-trained conditions.
+We leverage this geometry to shape the latent feature space in the challenging multi-label FL setting, where client data often deviates from the global distribution.
+An overview of our framework is shown in Figure 2.
+Label-Aware Disentanglement Module (LADM)
+Multi-label images may contain multiple object categories, making a single pooled representation suboptimal because it compresses multiple semantics into one vector.
+Inspired by recent advances in multi-label learning (Liu et al.,, 2021; Ridnik et al.,, 2023; Xu et al.,, 2023), we incorporate an attention mechanism to extract class-specific information from the image-level representation, thereby enhancing the discriminative power of each label.
+[…]
+$\mathbf{h}_{ic}\;=\;\operatorname{MultiHeadAttn}(\mathbf{m}_{c},\;\mathbf{f}_{i}),\quad c\in\mathcal{C}$ (3)
+and stacks them into
+$\mathbf{H}_{i}\;=\;\begin{bmatrix}\mathbf{h}_{i1}^{\top}\\[-2.0pt] \vdots\\[-2.0pt] \mathbf{h}_{iC}^{\top}\end{bmatrix}\in\mathbb{R}^{C\times d}.$ (4)
+Sharing the fixed query matrix $\mathbf{M}$ across clients mitigates aggregation conflicts and encourages globally aligned, disentangled features.
 Neural Collapse-Inspired Feature Alignment
-After disentangling image-level features into class-specific representations, each feature captures a distinct semantic concept.
-This enables the effective adaptation of established representation learning techniques from the multi-class setting to the multi-label context.
-Rather than allowing each local model to independently learn its own classification boundaries, which can be hindered by label imbalance or missing classes, we fix the classifier using a predefined simplex ETF matrix.
-This design enforces maximal angular separation between class embeddings, enhancing the model’s ability to distinguish between classes regardless of the local label distribution.
-By reducing the influence of local label co-occurrence, this approach also helps mitigate divergence between local and global data distributions.
-Once class-specific features are extracted via LADM, each feature is projected onto its corresponding ETF classifier weight to compute a class-specific logit.
-These logits are then passed through a sigmoid activation to obtain binary predictions.
-The model is trained using binary cross-entropy (BCE) loss.
-The process is formalised as follows:
-$\hat{y}_{i}^{c}=\text{Sigmoid}(\mathbf{h}^{c}_{i}\cdot\mathbf{m}_{c})$ (4)
-$\mathcal{L}_{\text{BCE}}=-\frac{1}{C}\frac{1}{N}\sum_{c=1}^{C}\sum_{i=1}^{N}\left[y_{i}^{c}\log(\hat{y}_{i}^{c})+(1-y_{i}^{c})\log(1-\hat{y}_{i}^{c})\right]$ (5)
-where $\mathbf{h}_{i}^{c}\in\mathbb{R}^{d}$ denotes the class-specific feature for class $c$ extracted by LADM for the ii‑th sample, and $\mathbf{m}_{c}\in\mathbb{R}^{d}$ represents the fixed ETF classifier weight corresponding to class $c$.
-$C$ is the total number of classes, and $N$ denotes the number of training samples available on the client.
+After disentangling image-level features into class-wise representations with LADM, each sample $i\in[N_{k}]$ yields a feature $\mathbf{h}_{ic}\in\mathbb{R}^{d}$ for every class $c\in\mathcal{C}$.
+To avoid client-specific classifier drift caused by label imbalance or missing classes, we fix the classifier to the global simplex ETF used as the LADM query matrix, $\mathbf{M}=[\mathbf{m}_{c}]_{c=1}^{C}\in\mathbb{R}^{d\times C}$.
+Given $\mathbf{h}_{ic}$ and its corresponding prototype $\mathbf{m}_{c}$, we compute a class logit and pass it through the sigmoid to obtain a binary prediction:
+$\hat{y}_{ic}\;=\;\sigma\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{c}\big),$ (5)
+and train with a binary cross-entropy (BCE) objective
+$\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\frac{1}{C}\sum_{i=1}^{N}\sum_{c=1}^{C}\Big[y_{ic}\,\log\hat{y}_{ic}+(1-y_{ic})\,\log(1-\hat{y}_{ic})\Big].$ (6)
+Here, $y_{ic}\in\{0,1\}$ is the ground-truth multi-label indicator, $\sigma(\cdot)$ denotes the sigmoid.
+This fixed classifier improves class separability and reduces divergence between local and global label distributions.
 Reducing Noise and Improving Clustering
-After extracting class-wise features using LADM, each sample is associated with multiple features, corresponding to the total number of classes.
-Based on the ground-truth labels, a subset of these features is considered as positive, while the remaining ones are treated as negative.
-To promote more effective clustering in the latent feature space, we introduce two additional regularization terms.
-These terms are designed to suppress potential noise introduced by negative features and to encourage compact clustering of positive features belonging to the same class.
-Negative Feature Rejection Loss: The $\mathcal{L}_{\text{BCE}}$ loss encourages each class-wise feature to align with its corresponding ETF prototype if it corresponds to a ground-truth (positive) label, and to be pushed away from its prototype if it corresponds to a negative label.
-To further reduce prediction ambiguity, it is important to prevent negative features from aligning with any other class prototypes.
-To address this, we compute the similarity between negative features and all class-wise prototypes, applying a penalty to high similarity scores.
-This regularization helps suppress the influence of negative features on the final prediction and mitigates potential noise.
-The process is formalised as follows:
-$\hat{y}_{ij}^{-}=\text{Sigmoid}(\mathbf{h}^{-}_{i}\cdot\mathbf{m}_{j})$ (6)
-$\mathcal{L}_{\text{Neg}}=-\frac{1}{C}\frac{1}{N}\sum_{j=1}^{C}\sum_{i=1}^{N}\mathbb{I}(\hat{y}_{ij}^{-}>\tau)\cdot\log(1-\hat{y}_{ij}^{-})$ (7)
-where $h_{i}^{-}\in\mathbb{R}^{d}$ denotes the negative feature of sample $i$.
-The threshold $\tau$ is a predefined value used to filter out negative features with low similarity scores, ensuring that only confident negative predictions contribute to the loss.
-In our experiments, we empirically set $\tau=0.3$.
-Positive Feature Contrastive Loss: To encourage more compact and discriminative class-wise clustering in the latent feature space, we introduce a contrastive loss.
-This loss enforces positive features to remain close to their corresponding class prototypes while being pushed away from all other class prototypes.
-The formulation is:
-$\mathcal{L}_{\text{Pos}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{\mathbf{h}^{c}_{i}\cdot\mathbf{m}_{c}}}{\sum_{j=1}^{C}e^{\mathbf{h}^{c}_{i}\cdot\mathbf{m}_{j}}}$ (8)
-where $\mathbf{h}^{c}_{i}\in\mathbb{R}^{d}$ denotes the positive feature of sample $i$ for class $c$. $\mathbf{m}_{j}$ represents the predefined prototype for class $j$.
-With the two regularization terms, one designed to suppress noise from negative features and the other to encourage compact clustering of positive features, the total loss can be expressed as:
-$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{BCE}}+\lambda_{1}\mathcal{L}_{\text{Neg}}+\lambda_{2}\mathcal{L}_{\text{Pos}}$ (9)
-where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters used to balance the BCE loss and the two regularization terms.
+Each sample produces $C$ class-wise features $\{\mathbf{h}_{ic}\}_{c=1}^{C}$.
+Let the positive and negative class sets for sample $i$ be $\mathcal{C}_{i}^{+}=\{\,c\mid y_{ic}=1\,\}$ and $\mathcal{C}_{i}^{-}=\{\,c\mid y_{ic}=0\,\}$, respectively.
+We introduce two regularizers: one to suppress noise from negative features and another to encourage compact clustering of positive features.
+Negative Feature Rejection Loss.
+While $\mathcal{L}_{\text{BCE}}$ discourages alignment of a negative feature $\mathbf{h}_{ic}$ with its own prototype $\mathbf{m}_{c}$ when $y_{ic}=0$, it does not prevent $\mathbf{h}_{ic}$ from spuriously aligning with other prototypes.
+We therefore penalize high similarity between every negative feature and all non-self prototypes:
+$\hat{s}_{icr}\;=\;\sigma\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{r}\big),\quad c\in\mathcal{C}_{i}^{-},\;r\in\mathcal{C}\setminus\{c\},$ (7)
+$\mathcal{L}_{\text{Neg}}=-\,\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathcal{C}_{i}^{-}|}\sum_{c\in\mathcal{C}_{i}^{-}}\frac{1}{C-1}\sum_{\begin{subarray}{c}r=1\\ r\neq c\end{subarray}}^{C}\mathbb{I}\!\big(\hat{s}_{icr}>\tau\big)\,\log\!\big(1-\hat{s}_{icr}\big).$ (8)
+The indicator $\mathbb{I}(\cdot)$ filters out low-similarity pairs so that only confident negatives contribute. In our experiments, we set $\tau=0.3$.
+Positive Feature Contrastive Loss.
+To encourage compact and discriminative class-wise clustering in the latent feature space, we introduce a contrastive loss.
+This loss enforces each positive feature $\mathbf{h}_{ic}$ to be closer to its own prototype than to others via a prototype-softmax:
+$\mathcal{L}_{\text{Pos}}=-\,\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathcal{C}_{i}^{+}|}\sum_{c\in\mathcal{C}_{i}^{+}}\log\frac{\exp\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{c}\big)}{\sum_{r=1}^{C}\exp\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{r}\big)}.$ (9)
+Total Objective.
+The total loss can be expressed as:
+$\mathcal{L}_{\text{total}}\;=\;\mathcal{L}_{\text{BCE}}+\lambda_{1}\,\mathcal{L}_{\text{Neg}}+\lambda_{2}\,\mathcal{L}_{\text{Pos}}.$ (10)
+where $\lambda_{1},\lambda_{2}\geq 0$ balance the contributions of the regularizers.
+FedNCAlign-ML
+We propose FedNCAlign-ML, a framework that enhances the standard FL protocol with an NC-inspired local training strategy tailored for multi-label tasks and skewed local data distributions.
+In this framework, each client first encodes its input images into features, which are then disentangled into class-specific embeddings using LADM.
+These embeddings are aligned with a shared simplex ETF matrix that serves simultaneously as fixed classifier prototypes and as class-wise queries for LADM, thereby enforcing a consistent global geometry across clients despite label imbalance.
+To further improve representation quality, we incorporate two regularizers: (i) a negative rejection loss to suppress spurious alignments of irrelevant features, and (ii) a positive contrastive loss to encourage compact intra-class clustering and clear inter-class separation.
+The novelty of FedNCAlign-ML lies in this local training design, which integrates disentangled class-wise embeddings, ETF-based alignment, and geometry-aware regularization.
+The complete procedure is summarized in Algorithm 1.
+Algorithm 1 FedNCAlign-ML: Federated Neural Collapse Alignment for Multi-Label Learning
+Input: $K$ clients with datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$; initial global model $w_{0}$; predefined ETF matrix $\mathbf{M}$; learning rate $\eta$; local epochs $E$; communication rounds $T$.
+1:Server executes:
+2:Initialize $w\leftarrow w_{0}$
+3:for $t=0$ to T1T-1 do $\triangleright$ communication rounds
+4:  for each client $k\in\{1,\dots,K\}$ in parallel do
+5:   $w_{t+1}^{k}\leftarrow\textsc{ClientUpdate}(\mathcal{D}_{k},w_{t},\mathbf{M})$
+6:  end for
+7:  $w_{t+1}\leftarrow\textsc{ModelAggregation}(\{w_{t+1}^{k}\}_{k=1}^{K})$
+8:  Broadcast $w_{t+1}$ to all clients
+9:end for
+10:
+11:function ClientUpdate($\mathcal{D}_{k},w,\mathbf{M}$)
+12:  for $e=1$ to $E$ do $\triangleright$ local epochs
+13:   for each batch $(x,y)\subset\mathcal{D}_{k}$ do
+14:     $\mathbf{f}\leftarrow\textsc{FeatureExtractor}(w,x)$
+15:     $\mathbf{H}\leftarrow\textsc{LADM}(\mathbf{f},\mathbf{M})$
+16:     Compute prediction $\hat{y}_{c}=\sigma(\mathbf{h}_{c}^{\top}\mathbf{m}_{c})$ $\triangleright$ Eqs. 5
+[…]
+18:     Update $w\leftarrow w-\eta\nabla_{w}\mathcal{L}_{\text{total}}$
+19:   end for
+20:  end for
+21:  return $w$
+22:end function
 $\beta=0.5,\gamma=0.5$ $\beta=0.1,\gamma=0.5$
@@
 Overall, across a range of non-IID configurations with limited class availability per client, our method consistently outperforms existing FL approaches, demonstrating strong robustness to various label-skewed data distributions.
-Ablation Study
-To evaluate the effectiveness of each component in the proposed method, we conducted an ablation study on the multi-label DermaMNIST dataset, with the client data distribution configured using $\beta=0.1$ and $\gamma=0.71$.
-The experimental results are presented in Table 5.
-As shown in the table, incorporating LADM for class-specific feature extraction, together with the predefined ETF classifier to encourage class-wise clustering alignment across clients in the latent feature space, yields improvements of 0.43% in class-wise AUC and 7.34% in class-wise F1 score.
-Additional performance gains are achieved by introducing regularisation terms that enhance the model’s discriminative ability, resulting in further improvements of 3.31% in class-wise AUC and 3.43% in class-wise F1 score.
 (a) Learnable FC classifier
@@
 Observing from subfigure (a), without feature disentanglement (LADM), the model relies on undesired information, such as the number of labels per sample, for clustering.
+(a) FC classifier
+(b) LADM + fixed ETF classifier
+(c) LADM + ETF + Dual regularization
+Figure 4: Pair-wise cosine similarity between test data class-wise average feature prototypes.
 $\beta=0.5,\gamma=0.5$ $\beta=0.1,\gamma=0.5$
@@
 Table 4: Comparison of related FL methods with ours on the ChestX-ray14 dataset (Wang et al.,, 2017).
+Ablation Study
+To evaluate the effectiveness of each component in the proposed method, we conducted an ablation study on the multi-label DermaMNIST dataset, with the client data distribution configured using $\beta=0.1$ and $\gamma=0.71$.
+The experimental results are presented in Table 5.
+As shown in the table, incorporating LADM for class-specific feature extraction, together with the predefined ETF classifier to encourage class-wise clustering alignment across clients in the latent feature space, yields improvements of 0.43% in class-wise AUC and 7.34% in class-wise F1 score.
+Additional performance gains are achieved by introducing regularisation terms that enhance the model’s discriminative ability, resulting in further improvements of 3.31% in class-wise AUC and 3.43% in class-wise F1 score.
 ETF Clf LADM $\mathcal{L}_{\text{Neg}}$ $\mathcal{L}_{\text{Pos}}$ macro-AUC macro-F1 AUC F1
@@
 Table 6: Ablation study of the class-wise feature extraction block - LADM on the multi-label DermaMNIST dataset.
-(a) FC classifier
-(b) LADM + fixed ETF classifier
-(c) LADM + ETF + Dual regularization
-Figure 4: Pair-wise cosine similarity between test data class-wise average feature prototypes.
 We also investigated the class-wise query mechanism within the attention-based LADM module.
@@
 Appendix
-Pseudo Code for FedNCAlign-ML
-We present the pseudocode for our proposed FedNCAlign-ML in Algorithm 1.
-The input comprises a set of clients, each with its own local dataset, and a centralized server responsible for managing communication.
-The training process involves three iterative steps: local training on client devices, aggregation at the central server, and broadcasting the updated global model back to the clients.
-While we adopt the standard FL protocol for aggregation and broadcasting, the core contribution of FedNCAlign-ML lies in its specialized model design and the local training procedure tailored for multi-label label-skewed FL scenarios.
-During local training, each client first encodes its input data using a standard feature extractor to generate image-level representations.
-These representations are then passed to the proposed Label Aware Disentanglement Module (LADM), which produces disentangled, class-wise embeddings.
-By capturing distinct semantic concepts within each feature, these embeddings facilitate more effective and controllable clustering in the latent feature space.
-To further enhance consistent and discriminative feature alignment across clients, particularly under label-skewed conditions, we introduce a shared, predefined Equiangular Tight Frame (ETF) structure.
-This structure encourages class-independent clustering and aligns local features with a global class geometry, helping to mitigate overfitting to local data distributions.
-Additionally, two complementary regularization terms are introduced to enhance feature clustering in the latent feature space: a rejection loss to filter out noisy signals from negative features, and a contrastive loss to promote compact intra-class clustering and clear inter-class separation.
-Algorithm 1 Federated Neural Collapse Alignment for Multi-Label Learning (FedNCAlign-ML)
-Input: $K$ clients; local datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$; initial global parameters $w_{0}$; predefined ETF matrix $\mathbf{M}$; learning rate $\eta$; local epochs $E$.
-1:Server:
-2:Initialize $w\leftarrow w_{0}$
-3:for $t=0$ to $T$ do $\triangleright$ $t$-th communication round
-4:  for each client $k=1,\dots,K$ in parallel do
-5:   $w_{t+1}^{k}\leftarrow\textsc{ClientUpdate}(\mathcal{D}_{k},w_{t})$
-6:   Send the client model $w^{k}_{t+1}$ to the server
-7:  end for
-8:  $w_{t+1}\leftarrow\textsc{ModelAggregation}(\{w_{t+1}^{k}\}_{k=1}^{K})$
-9:  Broadcast the global model $w_{t+1}$ to all clients
-10:end for
-11:ClientUpdate($\mathcal{D}_{k},w$):
-12:Split $\mathcal{D}_{k}$ into batches $\{b\}$
-13:for $e=0$ to $E$ do $\triangleright$ $e$-th local epoch
-14:  for each batch $b$ in $\{b\}$ do
-15:   Extract inputs $(x,y)$ from $b$
-16:   $\mathbf{f}\leftarrow\text{FeatureExtractor}(w;x)$
-17:   $\mathbf{H}\leftarrow\text{LADM}(\mathbf{M},\mathbf{f})$ $\triangleright$ Eq. 3
-18:   Compute loss $\mathcal{L}_{\text{total}}(w;\mathbf{M},\mathbf{H})$ $\triangleright$ Eqs. 5789
-19:   $w\leftarrow w-\eta\nabla_{w}\mathcal{L}_{\text{total}}$
-20:  end for
-21:end for
-22:Return $w$
 Ablation Study (additional)

### v3

@@
 Abstract
-Federated Learning (FL) enables collaborative model training across distributed clients while preserving data privacy.
-However, the performance of deep learning often deteriorates in FL due to decentralized and heterogeneous data.
-This challenge is further amplified in multi-label scenarios, where data exhibit complex characteristics such as label co-occurrence, inter-label dependency, and discrepancies between local and global label relationships.
-While most existing FL research primarily focuses on single-label classification, many real-world applications, particularly in domains such as medical imaging, often involve multi-label settings.
-In this paper, we address this important yet underexplored scenario in FL, where clients hold multi-label data with skewed label distributions.
-Neural Collapse (NC) describes a geometric structure in the latent feature space where features of each class collapse to their class mean with vanishing intra-class variance, and the class means form a maximally separated configuration.
-Motivated by this theory, we propose a method to align feature distributions across clients and to learn high-quality, well-clustered representations.
-To make the NC-structure applicable to multi-label settings, where image-level features may contain multiple semantic concepts, we introduce a feature disentanglement module that extracts semantically specific features.
-The clustering of these disentangled class-wise features is guided by a predefined shared NC structure, which mitigates potential conflicts between client models due to diverse local data distributions.
-In addition, we design regularisation losses to encourage compact clustering in the latent feature space.
-Experiments conducted on four benchmark datasets across eight diverse settings demonstrate that our approach outperforms existing methods, validating its effectiveness in this challenging FL scenario.
-Introduction
-Figure 1: Within an FL system, clients may present distinct data distributions and label relationships.
-For example, Client 1 primarily serves patients with sun exposure-related skin lesions, whereas Client 2 mainly handles cases involving melanoma and vascular lesions.
-The proposed method aims to address multi-label, label-skewed FL scenarios by disentangling class-specific features and promoting structured feature clustering in the latent feature space.
-Federated Learning (FL) has emerged as a promising framework for analysing sensitive data, particularly in domains like healthcare imaging, where privacy and legal constraints prevent direct data sharing.
-However, most existing FL approaches are limited to basic multi-class classification tasks, failing to account for more realistic scenarios where multiple labels frequently co-occur within a single sample.
-For instance, many diseases are interrelated, and patients often present with multiple conditions or symptoms simultaneously, rather than a single isolated issue.
-This multi-label nature of real-world data is often overlooked.
-Another core challenge in FL is data heterogeneity across clients, as each client (e.g., hospital) may possess a uniquely skewed data distribution.
-This variation arises from differences in geography, population demographics, medical facilities, and clinical expertise.
-As a result, clients may exhibit label imbalance, including majority, minority, or even missing (vacant) classes.
-Under multi-label scenarios, such heterogeneity becomes even more pronounced, making collaboration across clients significantly more difficult.
-Figure 1 illustrates this scenario, demonstrating that hospitals may hold datasets containing various sets of dermatological conditions.
-In this paper, we explore the problem of multi-label FL under label-skewed conditions, where each client’s local data is both imbalanced and divergent from the global distribution.
-Without directly sharing local data, the goal is to collaboratively train a global model that performs well across all target classes, despite the presence of class imbalance and missing labels on certain clients.
-To tackle the unique challenges posed by multi-label, label-skewed FL, we propose a novel method called Federated Neural Collapse Alignment for Multi-Label Learning (FedNCAlign-ML).
-Our approach introduces a representation learning framework for multi-label FL, inspired by Neural Collapse (NC) theory (Papyan et al.,, 2020), to enhance feature alignment across clients with heterogeneous label distributions.
-Specifically, we first encode each input image using a standard feature extractor, producing an image-level representation potentially containing multiple semantic concepts.
-We then leverage attention mechanisms to extract class-specific representations from this image-level representation, generating disentangled, class-wise embeddings.
-This effectively decomposes the complex multi-label problem into multiple single-label problems.
-Next, motivated by the NC phenomenon, we incorporate a predefined Equiangular Tight Frame (ETF) structure to promote consistent and discriminative feature alignment across clients, even under non-IID data distributions.
+Federated Learning (FL) enables collaborative model training across distributed clients while preserving data privacy, yet it remains challenging as data distributions can be highly heterogeneous.
+These challenges are further amplified in multi-label scenarios, where data exhibit characteristics such as label co-occurrence, inter-label dependency, and discrepancies between local and global label relationships.
+While most existing FL studies focus on single-label classification, real-world applications, such as in medical imaging, involve multi-label data with highly skewed label distributions across clients.
+To address this important yet underexplored problem, we propose FedNCA-ML, a novel FL framework that aligns feature distributions across clients and learns discriminative, well-clustered representations inspired by Neural Collapse (NC) theory.
+NC describes an ideal latent-space geometry where each class’s features collapse to their mean, forming a maximally separated simplex.
+To extend this theory to multi-label settings, we introduce a feature disentanglement module that extracts class-specific representations.
+The clustering of these disentangled features is guided by a shared NC-inspired structure, mitigating conflicts among client models caused by heterogeneous local data.
+Furthermore, we design regularisation losses to encourage compact and consistent feature clustering in the latent space.
+Experiments on four benchmark datasets under eight FL settings demonstrate the effectiveness of the proposed method, achieving improvements of up to 3.92% in class-wise AUC and 4.93% in class-wise F1 score.
+1 Introduction
+Figure 1: (a): Local client data are heterogeneous, exhibiting distinct class imbalances.
+(b): Multi-label data further exacerbate this heterogeneity through diverse label correlations.
+(c): FedNCA-ML addresses multi-label label-skewed FL challenges by disentangling class-specific features and promoting Neural Collapse–inspired structured clustering in the latent space.
+Federated Learning (FL) has emerged as a promising framework for analysing sensitive data, particularly in domains such as medical imaging, where privacy regulations and legal constraints prohibit direct data sharing.
+However, most existing FL approaches focus on basic multi-class classification tasks, overlooking more realistic scenarios in which multiple labels often co-occur within a single sample.
+For instance, many diseases are interrelated, and patients frequently present with multiple concurrent conditions.
+A fundamental challenge in FL is data heterogeneity across clients.
+Each client (e.g., hospital) may have a uniquely skewed data distribution due to variations in geography, population demographics, medical facilities, and clinical expertise.
+Consequently, local datasets often exhibit label imbalance, including majority, minority, or even missing classes.
+In multi-label scenarios, such heterogeneity becomes even more pronounced.
+Within each client, certain labels may frequently co-occur, forming client-specific label correlations that differ across clients.
+These inconsistent label relationships further complicate collaborative model training, posing significant challenges to achieving global model convergence and generalization.
+Figure 1 illustrates this scenario, showing that different clients hold datasets with diverse data distributions and label dependencies.
+In this paper, we explore the problem of multi-label FL under label-skewed conditions, where each client’s local data is both imbalanced and diverges from the global distribution.
+Without directly sharing local data, the goal is to collaboratively train a global model that performs well across all target classes, despite class imbalance, missing labels on certain clients, and diverse label correlations across clients.
+To tackle these challenges, we propose a novel method called Federated Neural Collapse Alignment for Multi-Label Learning (FedNCA-ML).
+FedNCA-ML introduces a unified representation learning framework for multi-label FL, inspired by Neural Collapse (NC) theory (papyan2020prevalence), to achieve consistent and discriminative feature alignment across heterogeneous clients.
+Specifically, each input image is first encoded by a standard feature extractor to obtain an image-level representation containing multiple semantic concepts.
+An attention-based disentanglement module then derives class-specific representations, effectively decomposing the complex multi-label problem into multiple single-label subproblems.
+Guided by NC theory, we further integrate a predefined Equiangular Tight Frame (ETF) structure to enforce consistent and discriminative feature alignment across clients.
 The ETF structure facilitates class-independent clustering, encouraging local models to align with the global class geometry rather than overfitting to their local data distributions.
-Additionally, we introduce two complementary regularisation terms to enhance feature clustering in the latent feature space.
-A rejection loss filters out noisy signals from negative features, and a contrastive loss encourages compact intra-class clustering and clear inter-class separation.
+In addition, two complementary regularisation terms enhance feature compactness and robustness.
+A rejection loss filters out noisy negative features, while a contrastive loss encourages compact intra-class clustering and clear inter-class separation.
 The key contributions of this paper are as follows:
-  • We explore the important yet underexplored problem of multi-label FL under label-skewed conditions. To this end, we propose FedNCAlign-ML, a novel method inspired by Neural Collapse theory.
-  • FedNCAlign-ML effectively adapts established representation learning techniques from the multi-class to the multi-label setting, and tackles the challenges introduced by heterogeneous local data distributions, label co-occurrence, and inter-label dependency.
-  • To handle the presence of multiple semantic concepts within multi-label samples, we introduce an attention-based feature disentanglement block that extracts class-specific representations from image-level features.
-Additionally, we incorporate two regularisation terms, one to suppress noisy negative features and another to promote compact class-wise clustering.
-  • We evaluate the proposed FedNCAlign-ML method on four benchmark datasets under eight diverse settings, comparing it to state-of-the-art FL approaches.
-The results show that FedNCAlign-ML consistently outperforms existing methods, achieving improvements of up to 2.87% in class-wise AUC and 7.29% in class-wise F1 score.
-Related Work
-Heterogeneous Federated Learning
-The foundational FL algorithm, FedAvg (McMahan et al.,, 2017), performs iterative weighted averaging of locally trained model parameters.
-This process involves local training on client devices, aggregation at a central server, and broadcasting the updated global model back to clients.
-While FedAvg performs well under independently and identically distributed (IID) data, real-world scenarios often involve non-IID data distributions.
-These distributions are commonly observed as quantity skew, label distribution skew, and feature distribution skew, all of which present significant challenges for FL.
-In the context of multi-class classification, numerous methods have been proposed to address such data heterogeneity within FL settings.
-For instance, FedProx (Li et al.,, 2020) adds an L2 regularisation term to constrain local updates and encourage proximity to the global model.
-FedCurv (Shoham et al.,, 2019) uses Elastic Weight Consolidation (Kirkpatrick et al.,, 2017) to preserve important local parameters to reduce divergence across clients.
-SCAFFOLD (Karimireddy et al.,, 2020) introduces control variates to correct client drift.
-SphereFed (Dong et al.,, 2022) projects features onto a hypersphere using a fixed classifier to mitigate latent feature distribution discrepancies.
-FedVLS (Guo et al.,, 2025) applies global model distillation to alleviate forgetting of missing classes and regularisation to amplify minority class losses, thereby improving performance on underrepresented classes.
-Although these approaches have demonstrated effectiveness in the multi-class setting, their performance tends to deteriorate when directly applied to the more complex multi-label scenario, where each sample may be associated with multiple co-occurring labels.
-In this work, we focus on addressing the challenges of multi-label FL under quantity-skewed and label-skewed data distribution.
-Multi-Label Federated Learning
-Although multi-label classification is a fundamental task, it remains less explored than multi-class classification in image processing.
-In the context of FL, the multi-label setting introduces additional and distinct challenges compared to the multi-class scenario, including label co-occurrence, inter-label dependency, and divergence between local and global label relationships.
-FedLGT (Liu et al.,, 2024) addresses these challenges by focusing on modelling label correlations across clients.
-Building upon C-Tran (Lanchantin et al.,, 2021), it leverages frozen CLIP (Radford et al.,, 2021) text embeddings to maintain consistent label relationships and reduce divergence in label dependencies among clients.
-In contrast, our approach focuses on understanding the structure of the latent feature space and explicitly controlling and optimising it to enhance robustness under challenging label-skewed FL conditions.
-Figure 2: Framework for the proposed FedNCAlign‑ML method for addressing multi‑label, label‑skewed FL.
-First, an attention‑based Label‑Aware Disentanglement Module (LADM) extracts label‑specific features from image‑level representations.
-Then, a predefined ETF matrix serves as both the shared classifier and the source of class-wise query embeddings, enabling consistent local training across clients.
-Additionally, two regularisation terms are introduced to suppress noisy negative features and encourage compact intra-class clustering in the latent feature space.
-Neural Collapse-Inspired Methods
-Recent research has shown that during the terminal phase of training (TPT), a phenomenon known as Neural Collapse (NC) emerges (Papyan et al.,, 2020).
-TPT refers to the stage when overparameterized deep networks are trained on balanced multi-class classification datasets, and the model’s performance plateaus while the training loss remains non-zero and continues to decrease.
+  • We explore the important yet underexplored problem of multi-label FL under label-skewed conditions, and propose FedNCA-ML, a novel NC-inspired framework for representation alignment in heterogeneous multi-label settings.
+  • FedNCA-ML proposes an attention-based disentanglement module to extract class-specific representations, extending NC principles to multi-label learning and mitigating data heterogeneity and label dependency across clients.
+  • Two regularization terms are designed to suppress noisy negative features and encourage compact class-wise compactness, improving global model generalization under label-skewed FL.
+2 Related Work
+2.1 Heterogeneous Federated Learning
+The classical FL algorithm FedAvg (mcmahan2017communication), which iteratively performs weighted averaging of locally trained model parameters, remains the foundation of most FL methods.
+Its process involves local training on client devices, aggregation at a central server, and broadcasting of the updated global model back to all clients.
+While FedAvg performs well under independently and identically distributed (IID) data, real-world applications often involve non-IID data distributions, which significantly hinder FL performance.
+Such heterogeneity typically arises from quantity skew, label distribution skew, and feature distribution skew.
+In the context of multi-class classification, many approaches have been proposed to alleviate these challenges in FL (li2020federated; shoham2019overcoming; karimireddy2020scaffold; dong2022spherefed; guo2025exploring; saha2025f).
+A detailed discussion of these works is provided in the Appendix A.
+However, their effectiveness degrades when directly applied to multi-label scenarios, where each sample may contain multiple co-occurring labels.
+In this work, we focus on addressing multi-label FL under quantity-skewed and label-skewed data distributions.
+2.2 Multi-Label Federated Learning
+In FL, the multi-label setting poses unique challenges, including label co-occurrence, inter-label dependency, and divergence between local and global label relationships.
+FedMLP (sun2024fedmlp) addresses these issues by allowing each client to exchange both its local model and class-wise positive and negative prototypes with the server to perform pseudo-labeling.
+FedLGT (liu2024language) focuses on modelling label correlations across clients.
+Building on C-Tran (lanchantin2021general), it leverages frozen CLIP (radford2021learning) text embeddings to maintain consistent label relationships and reduce divergence in label dependencies among clients.
+In contrast, our approach focuses on structuring and optimizing the latent feature space, explicitly enhancing its organization to improve model robustness under label-skewed multi-label FL conditions.
+Figure 2: Overview of the proposed FedNCA-ML framework for multi-label label-skewed FL.
+Subfigure (a) shows the overall architecture, while Subfigures (b)–(d) illustrate the Label-Aware Disentanglement Module (LADM) and the regularization losses.
+The attention-based LADM extracts label-specific features from image-level features.
+A predefined ETF matrix acts as both the shared classifier and the source of class-wise query embeddings, ensuring consistent local training across clients.
+Two regularisation terms are further incorporated to suppress noisy negative features and promote compact intra-class clustering in the latent feature space.
+2.3 Neural Collapse-Inspired Methods
+Recent studies have shown that during the terminal phase of training (TPT), a phenomenon known as Neural Collapse (NC) emerges (papyan2020prevalence).
+TPT refers to the stage in which overparameterized deep networks, trained on balanced multi-class classification datasets, reach performance saturation while the training loss continues to decrease.
 NC describes a distinctive geometric structure in the latent feature space: final-layer features collapse to their corresponding within-class means, and these class-wise prototypes align to form the vertices of a simplex Equiangular Tight Frame (ETF).
-As this simple yet elegant mathematical structure provides an optimal linear separable state for classification, NC-inspired classifiers have been explored in various domains, including federated learning (Li et al., 2023b, ), incremental learning (Yang et al., 2023b, ; Wei et al.,, 2025), few-shot learning (Yang et al., 2023b, ), and long-tail recognition (Gao et al.,, 2024).
-Despite growing interest, most existing work has focused on multi-class classification, while the equally important challenge of multi-label classification remains underexplored.
-To address this gap, Li et al., 2023a investigates the presence of NC in multi-label settings.
-They find that network training on a balanced multi-label dataset also exhibits NC but in a generalised form.
-In particular, features corresponding to single-label instances still conform to the Simplex ETF structure, while features associated with multi-label instances tend to approximate scaled averages of the single-label prototypes.
-MLC-NC (Tao et al.,, 2025) uses NC to process the long-tailed multi-label classification task in a centralised setting.
-However, it does not consider the missing class condition, which frequently arises in label-skewed FL scenarios across individual clients.
-To the best of our knowledge, our paper is the first to explore multi-label FL from an NC perspective.
-It opens the possibility to solve the heterogeneous multi-label FL challenge by fundamentally avoiding interference due to various local data distributions, label co-occurrence, and inter-label dependency.
-Preliminaries
-Problem Formulation
+This simple yet elegant geometric structure represents an optimally separable state for classification, motivating the development of NC-inspired classifiers across various tasks (li2023no; yang2023neural; wei2025compress; yang2023neural; gao2024distribution).
+Despite growing interest, most existing works focus on multi-class classification, leaving the equally important problem of multi-label classification underexplored.
+To bridge this gap, li2023neural investigated NC in multi-label settings and observed that models trained on balanced multi-label datasets exhibit a generalized form of NC.
+In particular, features corresponding to single-label instances still conform to the Simplex ETF structure, whereas features of multi-label samples approximate scaled averages of single-label prototypes.
+MLC-NC (tao2025mlc) further leverages NC to address long-tailed multi-label classification in a centralised setting.
+To the best of our knowledge, our paper is the first to explore multi-label FL from an NC perspective, opening new opportunities to address heterogeneous multi-label FL by mitigating interference caused by diverse local data distributions, label co-occurrence, and inter-label dependencies.
+3 Preliminaries
+3.1 Problem Formulation
 We consider a multi-label, label-skewed FL setting with $K$ clients collaboratively training a shared global model.
@@
 Let $\mathcal{C}=\{1,\dots,C\}$ denote the global class index set.
-Due to label skew, client $k$ may only observe a subset $\mathcal{C}_{k}\subseteq\mathcal{C}$, and even for shared classes, the class-conditional distributions can differ across clients.
-The goal is to learn a single global model that accurately recognizes all classes in $\mathcal{C}$ without sharing raw data among clients.
-Neural Collapse (NC)
-In this section, we first introduce the structure of the simplex ETF, followed by a discussion of the key properties of Neural Collapse.
+Due to label skew, client $k$ only observes a subset $\mathcal{C}_{k}\subseteq\mathcal{C}$, and even for shared classes, the class-conditional distributions can differ across clients.
+The goal is to learn a single global model that accurately recognizes all classes in $\mathcal{C}$ without sharing any local client data.
+3.2 Neural Collapse (NC)
+In this section, we first introduce the structure of the simplex ETF, followed by a discussion of the key properties of NC.
 Simplex Equiangular Tight Frame (ETF).
@@
 At the end of training, over-parameterized networks exhibit the NC phenomenon. Empirically, the following regularities are observed:
-  • Variability Collapse ($\mathcal{NC}_{1}$). Within-class feature variance collapses: features of the same class concentrate at their class mean.
-  • Convergence to Simplex ETF ($\mathcal{NC}_{2}$). Class means arrange as the vertices of a simplex ETF, forming a maximally symmetric and equidistant configuration.
-  • Self-Duality ($\mathcal{NC}_{3}$). Upon appropriate rescaling, the final-layer classifier weights align with the class means, sharing the same simplex ETF geometry.
-Proposed Method
-The NC phenomenon suggests a promising structure for optimal feature clustering in the latent feature space under well-trained conditions.
-We leverage this geometry to shape the latent feature space in the challenging multi-label FL setting, where client data often deviates from the global distribution.
-An overview of our framework is shown in Figure 2.
-Label-Aware Disentanglement Module (LADM)
-Multi-label images may contain multiple object categories, making a single pooled representation suboptimal because it compresses multiple semantics into one vector.
-Inspired by recent advances in multi-label learning (Liu et al.,, 2021; Ridnik et al.,, 2023; Xu et al.,, 2023), we incorporate an attention mechanism to extract class-specific information from the image-level representation, thereby enhancing the discriminative power of each label.
-[…]
-$\mathbf{h}_{ic}\;=\;\operatorname{MultiHeadAttn}(\mathbf{m}_{c},\;\mathbf{f}_{i}),\quad c\in\mathcal{C}$ (3)
-and stacks them into
+  • Variability Collapse ($\mathcal{NC}_{1}$). Within-class feature variance collapses, causing features of the same class to concentrate around their class mean.
+  • Convergence to Simplex ETF ($\mathcal{NC}_{2}$). Class means arrange themselves as the vertices of a simplex ETF, forming a maximally symmetric and equidistant configuration.
+  • Self-Duality ($\mathcal{NC}_{3}$). Upon appropriate rescaling, the final-layer classifier weights align with the class means, exhibiting the same simplex ETF geometry.
+4 Proposed Method
+NC theory suggests a promising structure for optimal feature clustering in the latent space.
+Building on this insight, we propose FedNCA-ML, a framework that leverages NC geometry to organize and regularize feature representations in the multi-label FL setting.
+An overview of the proposed framework is presented in Figure 2.
+Our method employs an attention-based disentanglement module (Fig.
+2b) to extract class-specific features from image-level representations and integrates a predefined ETF matrix as both the shared classifier and the source of class-wise query embeddings, enabling consistent feature alignment across clients (Fig.
+2a).
+Additionally, two regularization terms (Fig.
+2c, Fig.
+2d) are introduced to suppress noisy negative features and enhance intra-class compactness in the latent space.
+4.1 Label-Aware Disentanglement Module
+In multi-label images, representing the entire image with a single pooled feature vector is suboptimal, as it compresses diverse semantics into one embedding.
+To address this limitation, we introduce a Label-Aware Disentanglement Module (LADM) that extracts class-specific representations from the image-level features.
+Our design is motivated by recent advances in multi-label learning (liu2021query2label; ridnik2023ml; xu2023label) and the conceptual similarity between multi-label classification and object detection.
+In object detection, multiple bounding boxes may overlap, implying that a single region can contribute to multiple object instances.
+Analogously, in multi-label classification, certain image regions can simultaneously provide evidence for several semantic categories.
+Building upon this intuition, LADM adopts a cross-attention mechanism inspired by DETR (carion2020end).
+In LADM, a set of class queries attends to a grid of image tokens to extract class-specific representations.
+Unlike DETR, where each query corresponds to an instance and is trained with box-level supervision, LADM uses a fixed query per class and learns only from image-level labels.
+Each query functions as a soft region selector, aggregating spatial evidence relevant to its class while capturing contextual dependencies across regions.
+This design produces a set of disentangled, class-aware feature vectors that can be directly used for per-class prediction.
+In the FL setting, each client trains solely on its own local dataset, which may differ significantly from others due to non-IID data distributions.
+Such heterogeneity can cause inconsistencies in feature extraction across clients.
+To maintain global consistency, LADM employs a shared query matrix across all clients, ensuring that each class is represented by the same query direction.
+This shared design aligns feature distributions among clients and mitigates conflicts during model aggregation.
+Formally, for client $k$, each sample $x_{i}^{k}\in\mathcal{D}_{k}$ ($i\in[N_{k}]$) is encoded by the backbone into a spatial feature map:
+$\mathbf{F}_{i}^{k}\in\mathbb{R}^{d\times H^{\prime}\times W^{\prime}},$ (3)
+where $d$ denotes the channel dimension and $H\times W$ represents the spatial resolution. The feature map is then flattened into a sequence of $S=HW$ tokens:
+$\mathbf{Z}_{i}^{k}=\operatorname{reshape}(\mathbf{F}_{i}^{k})^{\top}\in\mathbb{R}^{S\times d},$ (4)
+and, for notational convenience, we omit the client index in subsequent equations and denote the token sequence as $\mathbf{Z}_{i}$.
+To preserve spatial structure, we add a fixed 2D sine–cosine positional embedding (carion2020end) to the key and value projections:
+$\tilde{\mathbf{Z}}_{i}=\mathbf{Z}_{i}+\mathbf{P},\quad\mathbf{P}=\mathrm{PE}_{2\mathrm{D}}(H^{\prime},W^{\prime},d),$ (5)
+where $\mathrm{PE}_{2\mathrm{D}}(\cdot)$ denotes the fixed sine–cosine positional encoding.
+This embedding introduces spatial awareness without adding learnable parameters, thereby ensuring a consistent inductive bias across clients.
+LADM employs a 4-head cross-attention mechanism to obtain class-specific representations from the encoded image tokens. We define a shared simplex ETF matrix as
+$\mathbf{M}=[\mathbf{m}_{c}]_{c=1}^{C}\in\mathbb{R}^{d\times C},$ (6)
+where each column $\mathbf{m}_{c}$ serves both as a fixed classifier weight and as a class-specific query vector.
+Given the query $\mathbf{m}_{c}$ and the encoded image tokens $\tilde{\mathbf{Z}}_{i}$, LADM computes the class-specific feature using multi-head cross-attention:
+$\mathbf{h}_{ic}\;=\;\operatorname{MultiHeadAttn}\bigl(\mathbf{m}_{c}^{\top},\,\tilde{\mathbf{Z}}_{i},\,\tilde{\mathbf{Z}}_{i}\bigr),\quad c\in\mathcal{C}.$ (7)
+The resulting features are stacked to form the class-feature matrix:
 $\mathbf{H}_{i}\;=\;\begin{bmatrix}\mathbf{h}_{i1}^{\top}\\[-2.0pt] \vdots\\[-2.0pt] \mathbf{h}_{iC}^{\top}\end{bmatrix}\in\mathbb{R}^{C\times d}.$ (8)
-Sharing the fixed query matrix $\mathbf{M}$ across clients mitigates aggregation conflicts and encourages globally aligned, disentangled features.
-Neural Collapse-Inspired Feature Alignment
+4.2 Neural Collapse-Inspired Feature Alignment
 After disentangling image-level features into class-wise representations with LADM, each sample $i\in[N_{k}]$ yields a feature $\mathbf{h}_{ic}\in\mathbb{R}^{d}$ for every class $c\in\mathcal{C}$.
-To avoid client-specific classifier drift caused by label imbalance or missing classes, we fix the classifier to the global simplex ETF used as the LADM query matrix, $\mathbf{M}=[\mathbf{m}_{c}]_{c=1}^{C}\in\mathbb{R}^{d\times C}$.
-Given $\mathbf{h}_{ic}$ and its corresponding prototype $\mathbf{m}_{c}$, we compute a class logit and pass it through the sigmoid to obtain a binary prediction:
+To prevent client-specific classifier drift, we fix the classifier to the global simplex ETF, which also serves as the LADM query matrix.
+This shared classifier enhances class separability and mitigates model drift caused by class imbalance and missing labels.
+Given $\mathbf{h}_{ic}$ and its corresponding prototype $\mathbf{m}_{c}$, we compute the class logit and apply the sigmoid function to obtain a binary prediction:
 $\hat{y}_{ic}\;=\;\sigma\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{c}\big),$ (9)
-and train with a binary cross-entropy (BCE) objective
+and train with a binary cross-entropy (BCE) loss:
 $\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\frac{1}{C}\sum_{i=1}^{N}\sum_{c=1}^{C}\Big[y_{ic}\,\log\hat{y}_{ic}+(1-y_{ic})\,\log(1-\hat{y}_{ic})\Big].$ (10)
-Here, $y_{ic}\in\{0,1\}$ is the ground-truth multi-label indicator, $\sigma(\cdot)$ denotes the sigmoid.
-This fixed classifier improves class separability and reduces divergence between local and global label distributions.
-Reducing Noise and Improving Clustering
+Here, $y_{ic}\in\{0,1\}$ is the ground-truth multi-label indicator, and $\sigma(\cdot)$ is the sigmoid function.
+4.3 Reducing Noise and Improving Clustering
 Each sample produces $C$ class-wise features $\{\mathbf{h}_{ic}\}_{c=1}^{C}$.
-Let the positive and negative class sets for sample $i$ be $\mathcal{C}_{i}^{+}=\{\,c\mid y_{ic}=1\,\}$ and $\mathcal{C}_{i}^{-}=\{\,c\mid y_{ic}=0\,\}$, respectively.
-We introduce two regularizers: one to suppress noise from negative features and another to encourage compact clustering of positive features.
+Let the positive and negative class sets for sample $i$ be defined as $\mathcal{C}_{i}^{+}=\{\,c\mid y_{ic}=1\,\}$ and $\mathcal{C}_{i}^{-}=\{\,c\mid y_{ic}=0\,\}$, respectively.
+We introduce two regularization terms: one to suppress noise from negative features, and another to promote compact clustering of positive features.
 Negative Feature Rejection Loss.
-While $\mathcal{L}_{\text{BCE}}$ discourages alignment of a negative feature $\mathbf{h}_{ic}$ with its own prototype $\mathbf{m}_{c}$ when $y_{ic}=0$, it does not prevent $\mathbf{h}_{ic}$ from spuriously aligning with other prototypes.
-We therefore penalize high similarity between every negative feature and all non-self prototypes:
+While $\mathcal{L}_{\text{BCE}}$ discourages alignment between a negative feature $\mathbf{h}_{ic}$ and its corresponding prototype $\mathbf{m}_{c}$ when $y_{ic}=0$, it does not prevent $\mathbf{h}_{ic}$ from spuriously aligning with other class prototypes.
+To address this, we introduce a penalty on high similarity between each negative feature and all non-self prototypes:
 $\hat{s}_{icr}\;=\;\sigma\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{r}\big),\quad c\in\mathcal{C}_{i}^{-},\;r\in\mathcal{C}\setminus\{c\},$ (11)
-$\mathcal{L}_{\text{Neg}}=-\,\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathcal{C}_{i}^{-}|}\sum_{c\in\mathcal{C}_{i}^{-}}\frac{1}{C-1}\sum_{\begin{subarray}{c}r=1\\ r\neq c\end{subarray}}^{C}\mathbb{I}\!\big(\hat{s}_{icr}>\tau\big)\,\log\!\big(1-\hat{s}_{icr}\big).$ (8)
-The indicator $\mathbb{I}(\cdot)$ filters out low-similarity pairs so that only confident negatives contribute. In our experiments, we set $\tau=0.3$.
+and define the negative feature rejection loss as
+$\begin{split}\mathcal{L}_{\text{Neg}}&=-\,\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathcal{C}_{i}^{-}|}\sum_{c\in\mathcal{C}_{i}^{-}}\\[4.0pt] &\quad\times\frac{1}{C-1}\sum_{\begin{subarray}{c}r=1\\ r\neq c\end{subarray}}^{C}\mathbb{I}\!\big(\hat{s}_{icr}>\tau\big)\,\log\!\big(1-\hat{s}_{icr}\big)\end{split}$ (12)
+The indicator $\mathbb{I}(\cdot)$ filters out low-similarity pairs, ensuring that only confident negatives contribute to the loss. In our experiments, we set $\tau=0.3$.
 Positive Feature Contrastive Loss.
 To encourage compact and discriminative class-wise clustering in the latent feature space, we introduce a contrastive loss.
-This loss enforces each positive feature $\mathbf{h}_{ic}$ to be closer to its own prototype than to others via a prototype-softmax:
+This loss drives each positive feature $\mathbf{h}_{ic}$ to be closer to its own prototype than to others through a prototype-based softmax:
 $\mathcal{L}_{\text{Pos}}=-\,\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathcal{C}_{i}^{+}|}\sum_{c\in\mathcal{C}_{i}^{+}}\log\frac{\exp\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{c}\big)}{\sum_{r=1}^{C}\exp\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{r}\big)}.$ (13)
-Total Objective.
-The total loss can be expressed as:
+Total Objective. The total training loss is defined as:
 $\mathcal{L}_{\text{total}}\;=\;\mathcal{L}_{\text{BCE}}+\lambda_{1}\,\mathcal{L}_{\text{Neg}}+\lambda_{2}\,\mathcal{L}_{\text{Pos}}.$ (14)
-where $\lambda_{1},\lambda_{2}\geq 0$ balance the contributions of the regularizers.
-FedNCAlign-ML
-We propose FedNCAlign-ML, a framework that enhances the standard FL protocol with an NC-inspired local training strategy tailored for multi-label tasks and skewed local data distributions.
-In this framework, each client first encodes its input images into features, which are then disentangled into class-specific embeddings using LADM.
-These embeddings are aligned with a shared simplex ETF matrix that serves simultaneously as fixed classifier prototypes and as class-wise queries for LADM, thereby enforcing a consistent global geometry across clients despite label imbalance.
-To further improve representation quality, we incorporate two regularizers: (i) a negative rejection loss to suppress spurious alignments of irrelevant features, and (ii) a positive contrastive loss to encourage compact intra-class clustering and clear inter-class separation.
-The novelty of FedNCAlign-ML lies in this local training design, which integrates disentangled class-wise embeddings, ETF-based alignment, and geometry-aware regularization.
-The complete procedure is summarized in Algorithm 1.
-Algorithm 1 FedNCAlign-ML: Federated Neural Collapse Alignment for Multi-Label Learning
+where $\lambda_{1},\lambda_{2}\geq 0$ are weighting coefficients that balance the contributions of the regularization terms.
+4.4 FedNCA-ML
+In summary, we propose FedNCA-ML, a framework that enhances the standard FL protocol with an NC-inspired local training strategy for multi-label tasks under skewed client data distributions.
+Each client encodes images into features that are disentangled into class-specific embeddings via LADM, which are then aligned with a shared simplex ETF matrix serving as both fixed classifier prototypes and class-wise queries.
+This alignment enforces a consistent global geometry across clients despite label imbalance.
+Two regularizers further refine representation quality: a negative rejection loss to suppress spurious alignments, and a positive contrastive loss to promote compact intra-class clustering and clear inter-class separation.
+This geometry-aware training framework unifies disentangled embeddings, ETF-based alignment, and structure-preserving regularization.
+The overall procedure is summarized in Algorithm 1.
+Algorithm 1 FedNCA-ML
 Input: $K$ clients with datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$; initial global model $w_{0}$; predefined ETF matrix $\mathbf{M}$; learning rate $\eta$; local epochs $E$; communication rounds $T$.
@@
 6:  end for
-7:  $w_{t+1}\leftarrow\textsc{ModelAggregation}(\{w_{t+1}^{k}\}_{k=1}^{K})$
+7:  $w_{t+1}\leftarrow\frac{1}{K}\sum_{k=1}^{K}w_{t+1}^{k}$ $\triangleright$ model aggregation
 8:  Broadcast $w_{t+1}$ to all clients
@@
 11:function ClientUpdate($\mathcal{D}_{k},w,\mathbf{M}$)
-12:  for $e=1$ to $E$ do $\triangleright$ local epochs
+12:  for $e=0$ to E1E-1 do $\triangleright$ local epochs
 13:   for each batch $(x,y)\subset\mathcal{D}_{k}$ do
 14:     $\mathbf{F}\leftarrow\textsc{FeatureExtractor}(w,x)$
-15:     $\mathbf{H}\leftarrow\textsc{LADM}(\mathbf{f},\mathbf{M})$
+15:     $\mathbf{H}\leftarrow\textsc{LADM}(\mathbf{F},\mathbf{M})$ $\triangleright$ Eqs. 345678
 16:     Compute prediction $\hat{y}_{c}=\sigma(\mathbf{h}_{c}^{\top}\mathbf{m}_{c})$ $\triangleright$ Eqs. 9
@@
 22:end function
-$\beta=0.5,\gamma=0.5$ $\beta=0.1,\gamma=0.5$
-macro-AUC macro-F1 AUC F1 macro-AUC macro-F1 AUC F1
-Centralized (Upperbound) 91.17 61.40 91.04 62.22 91.17 61.40 91.04 62.22
-FedAvg (McMahan et al.,, 2017) 82.46 40.65 81.83 41.24 78.84 31.15 77.00 34.98
-FedCurv (Shoham et al.,, 2019) 82.19 39.36 81.81 39.61 79.03 32.31 77.97 36.13
-FedProx (Li et al.,, 2020) 82.23 39.37 81.68 39.27 79.40 31.01 76.69 35.73
-SCAFFOLD (Karimireddy et al.,, 2020) 82.01 38.41 81.49 38.83 79.29 32.38 77.74 35.90
-SphereFed (Dong et al.,, 2022) 84.85 47.58 85.04 47.29 83.27 39.52 79.45 40.77
-FedLGT (Liu et al.,, 2024) 84.24 45.45 84.16 45.99 81.44 35.95 81.49 38.78
+Table 1: Comparisons on multi-label CIFAR-10 (krizhevsky2009learning) with the class presence ratio ($\gamma$) set to 0.5 ($\leq$ 5 of 10 classes per client).
+Non-IID client distributions are simulated using the Dirichlet factor ($\beta$).
+Class-wise (macro) and instance-wise (micro) performance are reported.
+$\beta=0.5$ $\beta=0.1$
+macro-AUC macro-F1 micro-AUC micro-F1 macro-AUC macro-F1 micro-AUC micro-F1
+Centralized (Upperbound) 92.20 $\pm$ 0.36 61.30 $\pm$ 0.94 90.04 $\pm$ 0.65 62.02 $\pm$ 0.62 92.20 $\pm$ 0.36 61.30 $\pm$ 0.94 90.04 $\pm$ 0.65 62.02 $\pm$ 0.62
+FedAvg (mcmahan2017communication) 82.29 $\pm$ 0.46 39.47 $\pm$ 1.59 81.48 $\pm$ 0.78 40.26 $\pm$ 1.09 78.92 $\pm$ 0.39 31.17 $\pm$ 0.51 77.62 $\pm$ 0.24 35.07 $\pm$ 0.47
+FedCurv (shoham2019overcoming) 82.53 $\pm$ 0.30 39.96 $\pm$ 1.28 82.10 $\pm$ 0.29 40.34 $\pm$ 1.04 79.06 $\pm$ 0.51 31.00 $\pm$ 1.21 77.46 $\pm$ 0.47 35.03 $\pm$ 0.54
+FedProx (li2020federated) 82.45 $\pm$ 0.34 39.22 $\pm$ 0.74 81.77 $\pm$ 0.48 39.66 $\pm$ 0.53 78.82 $\pm$ 0.55 30.74 $\pm$ 0.47 77.40 $\pm$ 0.31 35.02 $\pm$ 0.27
+SCAFFOLD (karimireddy2020scaffold) 82.51 $\pm$ 0.44 39.98 $\pm$ 1.41 82.08 $\pm$ 0.53 40.26 $\pm$ 1.26 79.00 $\pm$ 0.23 31.38 $\pm$ 0.31 77.72 $\pm$ 0.15 35.54 $\pm$ 0.16
+SphereFed (dong2022spherefed) 83.63 $\pm$ 1.50 42.58 $\pm$ 3.20 83.50 $\pm$ 1.87 43.18 $\pm$ 2.34 80.62 $\pm$ 1.46 36.83 $\pm$ 1.39 78.37 $\pm$ 0.95 38.18 $\pm$ 1.40
+FedLGT (liu2024language) 83.52 $\pm$ 0.68 43.60 $\pm$ 1.68 83.36 $\pm$ 0.71 44.03 $\pm$ 1.71 80.54 $\pm$ 0.43 36.30 $\pm$ 0.82 80.65 $\pm$ 0.68 39.24 $\pm$ 0.71
 […]
-Table 1: Comparisons on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009).
-Non-IID data distributions across clients are simulated using the Dirichlet concentration parameter ($\beta$) and the class-missing ratio ($\gamma$).
-$\beta=0.5,\gamma=0.71$ $\beta=0.1,\gamma=0.71$
-macro-AUC macro-F1 AUC F1 macro-AUC macro-F1 AUC F1
-Centralized (Upperbound) 91.86 63.49 94.52 73.30 91.86 63.49 94.52 73.30
-FedAvg (McMahan et al.,, 2017) 90.18 53.96 92.68 67.53 83.95 40.69 86.68 63.03
-FedCurv (Shoham et al.,, 2019) 90.31 56.17 92.64 68.11 86.20 42.19 89.54 64.76
-FedProx (Li et al.,, 2020) 90.16 55.85 92.44 67.95 83.73 42.98 87.96 63.50
-SCAFFOLD (Karimireddy et al.,, 2020) 89.56 54.88 92.13 67.59 85.33 43.60 89.41 64.75
-SphereFed (Dong et al.,, 2022) 85.36 44.25 89.46 53.40 83.37 41.90 88.92 50.94
-FedLGT (Liu et al.,, 2024) 87.33 56.24 90.84 68.25 84.80 44.09 88.70 57.35
+Table 2: Comparisons on multi-label DermaMNIST (yang2023medmnist) with the class presence ratio ($\gamma$) set to 0.71 ($\leq$ 5 of 7 classes per client).
+$\beta=0.5$ $\beta=0.1$
+macro-AUC macro-F1 micro-AUC micro-F1 macro-AUC macro-F1 micro-AUC micro-F1
+Centralized (Upperbound) 91.83 $\pm$ 0.40 64.38 $\pm$ 1.01 94.48 $\pm$ 0.25 74.12 $\pm$ 0.70 91.83 $\pm$ 0.40 64.38 $\pm$ 1.01 94.48 $\pm$ 0.25 74.12 $\pm$ 0.70
+FedAvg (mcmahan2017communication) 89.72 $\pm$ 0.29 54.88 $\pm$ 0.97 92.51 $\pm$ 0.36 68.29 $\pm$ 0.45 83.88 $\pm$ 1.21 42.79 $\pm$ 2.03 87.23 $\pm$ 0.86 62.81 $\pm$ 0.83
+FedCurv (shoham2019overcoming) 89.48 $\pm$ 0.18 54.92 $\pm$ 1.11 92.26 $\pm$ 0.36 67.97 $\pm$ 0.92 85.53 $\pm$ 1.37 43.45 $\pm$ 1.20 88.87 $\pm$ 0.50 64.37 $\pm$ 0.26
+FedProx (li2020federated) 89.69 $\pm$ 0.31 55.78 $\pm$ 0.94 92.06 $\pm$ 0.46 68.01 $\pm$ 1.46 84.45 $\pm$ 1.34 43.54 $\pm$ 1.35 88.28 $\pm$ 1.07 63.25 $\pm$ 0.63
+SCAFFOLD (karimireddy2020scaffold) 89.72 $\pm$ 0.49 55.85 $\pm$ 0.68 92.45 $\pm$ 0.32 68.09 $\pm$ 0.59 84.59 $\pm$ 1.03 43.20 $\pm$ 1.00 88.91 $\pm$ 0.90 63.82 $\pm$ 1.16
+SphereFed (dong2022spherefed) 85.09 $\pm$ 0.87 43.41 $\pm$ 1.73 89.65 $\pm$ 0.55 65.24 $\pm$ 0.58 81.78 $\pm$ 1.14 40.90 $\pm$ 1.79 86.59 $\pm$ 1.87 54.94 $\pm$ 1.12
+FedLGT (liu2024language) 87.63 $\pm$ 0.71 55.82 $\pm$ 1.34 91.19 $\pm$ 0.51 67.93 $\pm$ 0.45 84.91 $\pm$ 0.79 45.61 $\pm$ 1.16 89.42 $\pm$ 0.65 61.15 $\pm$ 2.43
 […]
-Table 2: Comparisons on the multi-label DermaMNIST dataset (Yang et al., 2023a, ).
-$\beta=0.05,\gamma=0.5$ $\beta=0.01,\gamma=0.5$
-macro-AUC macro-F1 AUC F1 macro-AUC macro-F1 AUC F1
-Centralized (Upperbound) 95.66 74.77 96.22 77.15 95.66 74.77 96.22 77.15
-FedAvg (McMahan et al.,, 2017) 93.72 57.00 94.20 64.62 93.29 49.16 93.90 61.08
-FedCurv (Shoham et al.,, 2019) 93.69 57.32 94.40 64.74 93.33 48.93 93.87 61.35
-FedProx (Li et al.,, 2020) 93.70 57.08 94.26 64.46 93.05 48.99 93.74 61.16
-SCAFFOLD (Karimireddy et al.,, 2020) 93.07 58.20 93.97 65.86 93.44 50.89 93.93 62.10
-SphereFed (Dong et al.,, 2022) 85.72 33.71 87.38 38.19 86.85 32.44 87.94 36.18
-FedLGT (Liu et al.,, 2024) 91.07 61.40 91.39 65.88 91.20 56.59 91.20 64.58
+Table 3: Comparisons on PASCAL VOC (everingham2010pascal) with the class presence ratio ($\gamma$) set to 0.5 ($\leq$ 10 of 20 classes per client).
+$\beta=0.05$ $\beta=0.01$
+macro-AUC macro-F1 micro-AUC micro-F1 macro-AUC macro-F1 micro-AUC micro-F1
+Centralized (Upperbound) 95.48 $\pm$ 0.11 74.61 $\pm$ 0.10 96.11 $\pm$ 0.12 76.94 $\pm$ 0.09 95.48 $\pm$ 0.11 74.61 $\pm$ 0.10 96.11 $\pm$ 0.12 76.94 $\pm$ 0.09
+FedAvg (mcmahan2017communication) 93.54 $\pm$ 0.23 57.57 $\pm$ 0.76 94.13 $\pm$ 0.33 64.67 $\pm$ 0.35 93.31 $\pm$ 0.11 47.73 $\pm$ 1.03 93.05 $\pm$ 0.19 60.46 $\pm$ 0.43
+FedCurv (shoham2019overcoming) 93.53 $\pm$ 0.33 57.36 $\pm$ 0.41 94.10 $\pm$ 0.20 64.60 $\pm$ 0.26 93.13 $\pm$ 0.29 48.54 $\pm$ 0.68 93.81 $\pm$ 0.09 60.49 $\pm$ 1.16
+FedProx (li2020federated) 93.55 $\pm$ 0.20 57.04 $\pm$ 0.24 94.04 $\pm$ 0.17 64.43 $\pm$ 0.29 93.14 $\pm$ 0.25 49.49 $\pm$ 0.68 93.80 $\pm$ 0.17 62.18 $\pm$ 0.51
+SCAFFOLD (karimireddy2020scaffold) 93.17 $\pm$ 0.20 57.44 $\pm$ 0.33 94.03 $\pm$ 0.11 64.95 $\pm$ 0.21 93.41 $\pm$ 0.17 49.64 $\pm$ 0.77 93.60 $\pm$ 0.21 61.44 $\pm$ 0.56
+SphereFed (dong2022spherefed) 83.72 $\pm$ 1.82 33.49 $\pm$ 1.15 84.38 $\pm$ 1.10 38.19 $\pm$ 1.17 84.25 $\pm$ 2.44 32.44 $\pm$ 0.21 85.94 $\pm$ 3.67 35.18 $\pm$ 1.30
+FedLGT (liu2024language) 91.93 $\pm$ 0.42 62.11 $\pm$ 0.44 91.75 $\pm$ 0.42 67.58 $\pm$ 0.45 91.25 $\pm$ 0.27 56.53 $\pm$ 2.42 91.46 $\pm$ 0.41 63.51 $\pm$ 1.73
 […]
-Table 3: Comparison of related FL methods with ours on the multi-label PASCAL VOC 2012 dataset (Everingham et al.,, 2010).
-Experiment
-Dataset and Evaluation Metric
+5 Experiments
+5.1 Dataset and Evaluation Metric
 Datasets: We conduct experiments on both general computer vision (CV) datasets and medical imaging datasets to evaluate the effectiveness and real-world applicability of the proposed method.
-For general CV tasks, we use CIFAR-10 (Krizhevsky et al.,, 2009) and PASCAL VOC (Everingham et al.,, 2010).
-Since CIFAR-10 is originally a multi-class dataset, we adopt the method used by Li et al., 2023a to transform it into a multi-label dataset.
+For general CV tasks, we use CIFAR-10 (krizhevsky2009learning) and PASCAL VOC (everingham2010pascal).
+Since CIFAR-10 is originally a multi-class dataset, we adopt the method used by li2023neural to transform it into a multi-label dataset.
 This is done by merging multiple images into a single composite image, with the combined labels serving as the new ground truth.
-For medical imaging tasks, we use DermaMNIST (Yang et al., 2023a, ) and ChestX-ray14 (Wang et al.,, 2017).
+For medical imaging tasks, we use DermaMNIST (yang2023medmnist) and ChestX-ray14 (wang2017chestxray).
 DermaMNIST, which originally contains 7 skin disease categories in a multi-class format, is converted to a multi-label dataset using the same strategy applied to CIFAR-10.
 ChestX-ray14 is naturally a multi-label dataset, comprising 14 thoracic disease categories and an additional “No Finding” label.
-Since a significant portion of the dataset, approximately four-sevenths of the training data, are “No Finding” samples (negative cases with all-zero labels), we distribute these samples evenly across all clients.
+Since a significant portion of the dataset, 57% of the training data, is “No Finding” samples (negative cases with all-zero labels), we distribute these samples evenly across all clients.
 This setup mimics a realistic clinical scenario in which healthy cases are prevalent, while disease cases are relatively rare and unevenly distributed across clients.
-Detailed information about the datasets and local data distributions under various experimental settings is provided in the Appendix.
-Evaluation Metric: Given our focus on label-skewed data distributions, where certain classes may be missing or underrepresented, we report both overall (micro) and class-wise (macro) performance metrics.
-Following standard practices across different datasets, we report AUC and F1 scores for CIFAR-10, PASCAL VOC, and DermaMNIST, and AUC for ChestX-ray14.
-Together, these metrics offer a balanced and comprehensive evaluation by capturing both overall performance and class-specific behaviour.
-Task Setup and Implementation Details
-Task Setup: We simulate the FL system with 10 clients and adopt a full client participation strategy.
-To reflect the standard heterogeneous FL setting, we introduce non-IID data distributions by partitioning the data using a Dirichlet distribution, parameterised by the concentration factor $\beta$.
-To further emulate label-skewed scenarios, we define a class presence ratio $\gamma$, which limits the number of classes available to each client and simulates cases where certain classes are entirely absent from local datasets.
-Implementation Details: For all experiments, we use ResNet-18 (He et al.,, 2016) as the feature extractor.
-Each FL global model is trained over 100 communication rounds, with one epoch of local training per round.
-The final model is selected based on its best performance on the validation set.
-We use a batch size of 32 and initialise the learning rate at $1\times 10^{-4}$, using the AdamW optimiser with a weight decay of 0.01.
-Models for the CIFAR-10 dataset are trained from scratch, while models for other datasets are initialised with ImageNet-pretrained weights.
-During training, the regularisation hyperparameter $\lambda_{1}$ is empirically set to either 1 or 0.01, depending on the background complexity of the dataset.
-For datasets with complex backgrounds, where disentangling class-specific features is more challenging, a smaller value ($\lambda_{1}=0.01$) is used to reduce potential interference with positive feature learning.
-The other regularisation hyperparameter, $\lambda_{2}$, is fixed at 1 across all experiments.
-Performance Comparison
-To evaluate the effectiveness of the proposed method, we compare it with other methods across four datasets, as shown in Tables 3, 3, 3, and 4.
-As observed in these tables, our method consistently achieves the best performance in most cases across different datasets.
-The improvement is particularly notable on naturally long-tailed datasets such as DermaMNIST and ChestX-ray14, where our method yields substantial gains in class-wise performance, indicating its ability to produce more balanced predictions.
-Specifically, under the $\beta=0.1$ setting on DermaMNIST, our method outperforms the second-best approach by 1.49% in class-wise AUC and 7.29% in class-wise F1 score.
-For PASCAL VOC, when each client is limited to a maximum of half the total classes, our method achieves improvements of 2.94% and 5.49% in class-wise F1 score under the $\beta=0.05$ and $\beta=0.01$ settings, respectively.
-For ChestX-ray14 under the $\beta=0.5$ and $\beta=0.1$ settings, our method improves class-wise AUC by 0.64% and 1.31%, respectively.
-Overall, across a range of non-IID configurations with limited class availability per client, our method consistently outperforms existing FL approaches, demonstrating strong robustness to various label-skewed data distributions.
+Detailed information about the datasets and local data distributions under various experimental settings is provided in the Appendix C.
+Evaluation Metric: Given our focus on label-skewed data distributions, we report both instance-wise (micro) and class-wise (macro) performance metrics.
+The macro metric provides a balanced evaluation across classes, mitigating bias toward frequent categories.
+Following standard practice, we report AUC and F1 scores for CIFAR-10, PASCAL VOC, and DermaMNIST, and AUC for ChestX-ray14.
+Together, these metrics comprehensively evaluate overall performance and class-level behaviour.
 (a) Learnable FC classifier
@@
 Observing from subfigure (a), without feature disentanglement (LADM), the model relies on undesired information, such as the number of labels per sample, for clustering.
-(a) FC classifier
-(b) LADM + fixed ETF classifier
-(c) LADM + ETF + Dual regularization
-Figure 4: Pair-wise cosine similarity between test data class-wise average feature prototypes.
-$\beta=0.5,\gamma=0.5$ $\beta=0.1,\gamma=0.5$
-macro-AUC AUC macro-AUC AUC
-Centralized 71.44 79.62 71.44 79.62
-FedAvg 69.07 73.28 69.84 78.00
-FedCurv 68.12 72.15 69.46 77.63
-FedProx 69.23 72.45 69.23 77.69
-SCAFFOLD 69.02 72.21 66.71 78.69
-SphereFed 57.48 70.26 61.80 74.59
-FedLGT 69.76 72.41 70.02 77.93
+5.2 Task Setup and Implementation Details
+Task Setup: We simulate the FL system with 10 clients and adopt a full client participation strategy.
+To reflect the standard heterogeneous FL setting, we introduce non-IID data distributions by partitioning the data using a Dirichlet distribution, parameterised by the concentration factor $\beta$.
+To further emulate label-skewed scenarios, we define a class presence ratio $\gamma$, which limits the number of classes available to each client and simulates cases where certain classes are entirely absent from local datasets.
+Implementation Details: All experiments adopt ResNet-18 (he2016deep) as the feature extractor.
+Each FL global model is trained for 100 communication rounds, with one local epoch per round, and the final model is selected based on its best validation performance.
+We use a batch size of 32, an initial learning rate of $1\times 10^{-4}$, and the AdamW optimizer with a weight decay of 0.01.
+Models on CIFAR-10 are trained from scratch, while models on other datasets are initialised with ImageNet-pretrained weights.
+The negative feature rejection regularization coefficient, $\lambda_{1}$, is empirically set to 1 for CIFAR-10 and 0.01 for other datasets.
+The positive feature contrastive loss coefficient, $\lambda_{2}$, is set to 1 across all experiments.
+To ensure fair comparison, all experiments are repeated three times with different random seeds, and the mean and standard deviation are reported.
+5.3 Performance Comparison
+To evaluate the effectiveness of the proposed method, we compare FedNCA-ML with state-of-the-art approaches across four datasets under eight label-skewed FL settings.
+The results, summarized in Tables 3, 3, 3, and 4, show that FedNCA-ML consistently achieves the best class-wise performance in most cases, demonstrating its strong capability for balanced and generalized predictions under heterogeneous FL conditions.
+Specifically, on multi-label CIFAR-10, under non-IID Dirichlet settings of $\beta=0.5$ ($\beta=0.1$) with a maximum of 5 out of 10 classes per client, FedNCA-ML surpasses the second-best approach by 3.92% (3.18%) in class-wise AUC and 4.57% (1.26%) in class-wise F1 score.
+On multi-label DermaMNIST, under $\beta=0.5$ ($\beta=0.1$) with up to 5 of 7 classes per client, it achieves improvements of 0.40% (0.46%) in AUC and 0.77% (4.93%) in F1 score.
+For PASCAL VOC, under more challenging non-IID settings of $\beta=0.05$ ($\beta=0.01$) with up to 10 of 20 classes per client, our method outperforms the second-best approach by 2.17% (4.55%) in class-wise F1 score.
+On ChestX-ray14, under non-IID Dirichlet settings of $\beta=0.5$ and $\beta=0.1$ with up to 7 of 14 disease classes per client, FedNCA-ML improves class-wise AUC by 0.69% and 1.21%, respectively.
+However, our method does not outperform others in overall AUC.
+ChestX-ray14 is highly imbalanced with 57% of training and 38% of testing samples labeled as “No Finding”.
+This severe imbalance leads many methods to overpredict the majority class, as reflected in a large gap between class-wise and overall AUC, indicating bias toward majority classes and degraded performance on minority (disease) classes.
+In medical diagnosis, false positives are generally more tolerable than false negatives, especially for rare diseases.
+The superior class-wise AUC and smaller gap between overall and class-wise AUC achieved by FedNCA-ML suggest more balanced predictions and better recognition of minority disease classes.
+Table 4: Comparisons on ChestX-ray14 (wang2017chestxray) with $\gamma$ set to 0.5.
+$\beta=0.5$ $\beta=0.1$
+macro-AUC micro-AUC macro-AUC micro-AUC
+Centralized 71.66 $\pm$ 0.12 79.54 $\pm$ 0.36 71.66 $\pm$ 0.12 79.54 $\pm$ 0.36
+FedAvg 69.02 $\pm$ 0.24 73.18 $\pm$ 0.15 69.05 $\pm$ 0.78 77.90 $\pm$ 0.20
+FedCurv 68.72 $\pm$ 1.19 72.15 $\pm$ 0.17 69.34 $\pm$ 0.48 77.36 $\pm$ 0.27
+FedProx 69.02 $\pm$ 0.21 72.04 $\pm$ 0.59 69.59 $\pm$ 0.23 77.43 $\pm$ 0.24
+SCAFFOLD 69.42 $\pm$ 0.39 72.49 $\pm$ 0.38 67.45 $\pm$ 0.54 77.90 $\pm$ 0.69
+SphereFed 58.35 $\pm$ 0.57 69.51 $\pm$ 0.56 61.96 $\pm$ 0.16 73.59 $\pm$ 1.02
+FedLGT 69.86 $\pm$ 0.76 72.27 $\pm$ 1.07 70.16 $\pm$ 0.37 77.67 $\pm$ 0.57
 […]
-Table 4: Comparison of related FL methods with ours on the ChestX-ray14 dataset (Wang et al.,, 2017).
-Ablation Study
-To evaluate the effectiveness of each component in the proposed method, we conducted an ablation study on the multi-label DermaMNIST dataset, with the client data distribution configured using $\beta=0.1$ and $\gamma=0.71$.
-The experimental results are presented in Table 5.
-As shown in the table, incorporating LADM for class-specific feature extraction, together with the predefined ETF classifier to encourage class-wise clustering alignment across clients in the latent feature space, yields improvements of 0.43% in class-wise AUC and 7.34% in class-wise F1 score.
-Additional performance gains are achieved by introducing regularisation terms that enhance the model’s discriminative ability, resulting in further improvements of 3.31% in class-wise AUC and 3.43% in class-wise F1 score.
-ETF Clf LADM $\mathcal{L}_{\text{Neg}}$ $\mathcal{L}_{\text{Pos}}$ macro-AUC macro-F1 AUC F1
+Table 5: Ablation study of the proposed method on the multi-label DermaMNIST dataset with $\beta=0.1$, $\gamma=0.71$.
+ETF Clf LADM $\mathcal{L}_{\text{Neg}}$ $\mathcal{L}_{\text{Pos}}$ macro-AUC macro-F1 micro-AUC micro-F1
 83.95 40.69 86.68 63.03
+83.61 33.35 87.26 61.48
 84.38 47.95 86.70 60.49
@@
 87.69 51.38 90.36 63.26
-Table 5: Ablation study of the proposed method on the multi-label DermaMNIST dataset with $\beta=0.1$, $\gamma=0.71$.
-query type query init macro-AUC macro-F1 AUC F1
+Table 6: Ablation study of the class-wise feature extraction block - LADM on the multi-label DermaMNIST dataset.
+query type query init macro-AUC macro-F1 micro-AUC micro-F1
 learnable random 82.94 47.94 86.33 57.37
@@
 fixed ETF 84.38 47.95 86.70 60.49
-Table 6: Ablation study of the class-wise feature extraction block - LADM on the multi-label DermaMNIST dataset.
-We also investigated the class-wise query mechanism within the attention-based LADM module.
-As reported in Table 6, fixed, well-designed class-wise queries consistently outperform learnable queries across most evaluation metrics.
-This outcome is intuitive in the context of label-skewed FL, where each client has a unique local data distribution.
-If the queries and classifier are learnable, the model may overfit to these local distributions, which can significantly deviate from the global distribution.
-Such overfitting introduces inconsistencies among client models and ultimately degrades the performance of the aggregated global model.
-To maintain consistent learning across clients with non-iid data, it is therefore preferable to use predefined, well-separated, fixed embeddings to guide the learning process rather than relying on learnable representations.
-To further analyse model behaviour, we present t-SNE visualisations of feature distributions from the test data in the latent feature space, generated under different model architectures and training strategies on the multi-label DermaMNIST dataset.
-The client data distribution is configured using $\beta=0.1$ and $\gamma=0.71$.
-As shown in Figure 3, when a conventional, learnable fully connected (FC) classifier is used, although the model achieves 83.95% class-wise AUC and 40.69% F1 score, the resulting feature representations exhibit poor clustering.
-Notably, the model appears to rely on undesired information.
-Specifically, it tends to group features based on the number of labels per sample, rather than focusing solely on semantic content.
-By incorporating LADM for extracting single-class features and employing a predefined ETF classifier to regulate feature distribution across clients, the model learns to focus more on semantic content for clustering.
-This leads to more disentangled, class-specific feature representations and improved clustering quality.
-Figure 4 also shows that the pairwise cosine similarity between class-wise average features is substantially reduced, indicating enhanced inter-class separability and stronger discriminative capability.
-Finally, with the addition of extra regularisation terms during training, the feature clustering becomes more compact and semantically coherent, and the similarity among class-wise average prototypes is further reduced.
-Conclusion
-This paper investigates multi-label label-skewed FL, a practically important yet underexplored problem setting.
-This task is particularly challenging due to the complexities inherent in multi-label data, combined with difficulties introduced by the federated system, including heterogeneous local data distributions and inconsistencies between local and global data distributions.
-Inspired by Neural Collapse theory, we propose a feature disentanglement module and incorporate a predefined ETF structure to facilitate the extraction and clustering of class-specific representations, while also aligning feature distributions across clients in the latent feature space.
-Furthermore, we incorporate specific regularization losses to suppress noisy negative features and enhance compact class-wise clustering.
-Experiments on four benchmark datasets across eight settings show that our method outperforms existing approaches, with gains of up to 2.87% in class-wise AUC and 7.29% in class-wise F1 score.
+(a) FC classifier
+(b) LADM + fixed ETF classifier
+(c) LADM + ETF + Dual regularization
+Figure 4: Pairwise cosine similarity of class-wise average feature prototypes.
+Incorporating LADM, ETF-based alignment, and structure-preserving regularization lowers inter-class similarity, reflecting enhanced separability and discrimination.
+Figure 5: Examples of attention maps from the multi-label DermaMNIST dataset.
+Each subfigure shows a sample image alongside the class-specific attention maps generated by the LADM module for the corresponding ground-truth labels.
+The LADM module captures class-specific features through its attention mechanism, where redder regions indicate areas of higher attention.
+5.4 Ablation Study
+To evaluate the effectiveness of each component in the proposed method, we conducted an ablation study on the multi-label DermaMNIST dataset.
+This dataset was selected because it exhibits a strong class imbalance and pronounced inter-class and intra-class variability, which arise from the diverse nature of dermatological conditions.
+In our experiments, the client data distribution was configured using $\beta=0.1$ and $\gamma=0.71$, which further intensifies the class imbalance and label heterogeneity.
+The results are summarized in Table 5.
+As shown in the table, solely using the predefined ETF classifier leads to degraded performance.
+This is because image-level features often contain multiple semantic cues, and enforcing direct clustering on such entangled representations confuses the model.
+In contrast, incorporating LADM for class-specific feature extraction, together with the predefined ETF classifier to promote class-wise clustering alignment across clients in the latent space, yields improvements of 0.43% in class-wise AUC and 7.34% in class-wise F1 score.
+Additional performance gains are achieved by introducing regularisation terms that enhance the model’s discriminative ability, resulting in further improvements of 3.31% in class-wise AUC and 3.43% in class-wise F1 score.
+We also investigated the attention mechanism within the LADM module.
+As shown in Table 6, fixed, well-designed class-wise queries consistently outperform learnable ones across most evaluation metrics.
+This result is intuitive under label-skewed FL, where clients have distinct local data distributions.
+When both the queries and the classifier are learnable, models tend to overfit local distributions that deviate from the global one, leading to inconsistencies among clients and degraded global performance after aggregation.
+In contrast, predefined, well-separated embeddings ensure consistent learning across clients with non-IID data.
+Figure 5 further illustrates that the LADM attention mechanism effectively focuses on semantically relevant image regions associated with the query class, thereby enabling discriminative class-specific feature extraction.
+To further analyse the model’s behaviour, we present t-SNE visualisations of the test data in the latent feature space (Figure 3) and the pairwise cosine similarity between class-wise average features (Figure 4), obtained under different model architectures and training strategies on the multi-label DermaMNIST dataset.
+As shown in Figure 3(a), when a conventional learnable fully connected (FC) classifier is used, the resulting feature representations exhibit poor clustering.
+Notably, the model appears to rely on undesired information, grouping features by the number of labels per sample rather than solely by semantic content.
+By incorporating LADM (Figure 3(b), 4(b)), which extracts single-class features, and further adding a predefined ETF classifier to regulate feature distribution across clients, the model learns to cluster features based on meaningful semantic attributes.
+This results in improved clustering quality, as evidenced by the substantially reduced pairwise cosine similarity between class-wise average features, indicating enhanced inter-class separability and stronger discriminative capability.
+Finally, incorporating additional regularization terms during training yields even more compact and semantically coherent feature clusters, with further reductions in prototype similarity (Figure 3(c), 4(c)).
+6 Conclusion
+This paper addresses the challenging yet underexplored problem of FL under multi-label and label-skewed conditions, where heterogeneous client data lead to severe class imbalance and inconsistent label correlations.
+To tackle these challenges, we propose FedNCA-ML, a Neural Collapse–inspired framework that enhances feature extraction and alignment across clients with non-IID data distributions.
+FedNCA-ML integrates a feature disentanglement module with a predefined ETF structure, extending Neural Collapse theory to the multi-label setting and facilitating consistent class-wise representation alignment among clients.
+Furthermore, two regularization losses are designed to suppress noisy negative features and enhance class-wise compactness in the latent space.
+Extensive experiments on four datasets under eight different FL settings demonstrate the effectiveness and robustness of the proposed method, yielding consistent improvement of up to 3.92% in class-wise AUC and 4.93% in class-wise F1 score.
 Acknowledgments
@@
 Appendix
-Ablation Study (additional)
+Appendix A Related Work (Additional)
+A.1 Heterogeneous Federated Learning
+The foundational Federated Learning (FL) algorithm, FedAvg (mcmahan2017communication), performs iterative weighted averaging of locally trained model parameters, and remains the underlying pipeline for many advanced FL methods.
+While FedAvg achieves strong performance under independently and identically distributed (IID) data, real-world scenarios often involve non-IID data distributions.
+Such heterogeneity commonly appears in the form of quantity skew, label-distribution skew, and feature-distribution skew, each of which poses substantial challenges for FL.
+In the multi-class classification setting, numerous methods have been proposed to address such data heterogeneity within FL settings.
+FedProx (li2020federated) introduces an L2 regularisation term to constrain local updates and encourage proximity to the global model.
+FedCurv (shoham2019overcoming) uses Elastic Weight Consolidation (kirkpatrick2017overcoming) to preserve important local parameters to reduce divergence across clients.
+SCAFFOLD (karimireddy2020scaffold) introduces control variates to correct client drift.
+SphereFed (dong2022spherefed) projects features onto a hypersphere using a fixed classifier to mitigate discrepancies in latent feature distributions.
+FedVLS (guo2025exploring) uses global model distillation to alleviate forgetting of missing classes and incorporates loss regularisation to amplify minority class signals, thereby improving performance on underrepresented classes.
+Although these approaches have demonstrated effectiveness in multi-class settings, their performance tends to deteriorate when directly applied to the more complex multi-label scenario, where each sample may be associated with multiple co-occurring labels.
+In this work, we focus on addressing the challenges of multi-label FL under quantity-skewed and label-skewed data distributions.
+Appendix B Experiment (Additional)
+B.1 Ablation Study (Additional)
 To complement the results reported in the main manuscript, we present an additional ablation study to evaluate the contribution of each component in our proposed method.
-The experiments are conducted on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009), with the client data distribution configured using a Dirichlet concentration parameter $\beta=0.5$ and a class-presenting ratio $\gamma=0.5$.
-The experimental results are summarized in Tables 7 and 8, and Figures 5 and 6.
+The experiments are conducted on the multi-label CIFAR-10 dataset (krizhevsky2009learning), where the client data distribution is configured using a Dirichlet concentration parameter of $\beta=0.5$ and a class-presence ratio of $\gamma=0.5$ (i.e., each client has access to at most 5 out of 10 classes).
+The experimental results are summarized in Tables 7 and 8, and Figures 6 and 7.
 As shown in Table 7, integrating LADM for class-specific feature extraction, along with the use of a predefined ETF classifier to encourage class-wise clustering alignment across clients in the latent feature space, leads to performance improvements of 3.38% in class-wise AUC and 3.99% in class-wise F1 score.
-Additional gains are achieved by introducing regularization terms that enhance the model’s discriminative capacity, resulting in a further class-wise increase of 1.88% in AUC and 4.58% in F1 score.
+Additional gains are achieved by introducing regularization terms that enhance the model’s discriminative capacity, resulting in a further class-wise increase of 1.88% in class-wise AUC and 4.58% in class-wise F1 score.
 Regarding LADM specifically, the results in Table 8 demonstrate that using fixed, well-designed class-wise queries is generally more effective than learnable queries across most evaluation metrics.
-To further analyze model behavior, we visualize the latent feature distributions of the test set using t-SNE under different architectural and training configurations.
-As shown in Figure 5, incorporating LADM for extracting single-class features, along with a predefined ETF classifier to regulate feature distribution across clients, enables the model to focus on semantic content rather than irrelevant factors such as the number of labels per sample.
-Furthermore, the addition of regularization terms during training leads to more compact and semantically coherent clusters, while also reducing the similarity among class-wise average prototypes.
-As shown in Figure 6, the pairwise cosine similarity between class-wise average features decreases with the inclusion of LADM and the ETF classifier, and is further reduced by the added regularization terms, indicating enhanced inter-class separability and stronger discriminative capability.
+Table 7: Ablation study of the proposed method on the multi-label CIFAR-10 dataset (krizhevsky2009learning). $\beta$ is set to 0.5 and $\gamma$ is set to 0.5.
 ETF Clf LADM $\mathcal{L}_{\text{Neg}}$ $\mathcal{L}_{\text{Pos}}$ macro-AUC macro-F1 AUC F1
 82.46 40.65 81.83 41.24
+83.44 15.29 81.75 16.77
 85.84 44.64 85.10 45.45
@@
 87.72 49.22 87.13 49.60
-Table 7: Ablation study of the proposed method on the multi-label CIFAR10 dataset. $\beta$ is set to 0.5 and $\gamma$ is set to 0.5.
+Table 8: Ablation study of the class-wise feature extraction block - LADM on the multi-label CIFAR-10 dataset.
 query type query init macro-AUC macro-F1 AUC F1
@@
 fixed ETF 85.84 44.64 85.10 45.45
-Table 8: Ablation study of the class-wise feature extraction block - LADM on the multi-label CIFAR10 dataset.
 (a) Learnable FC classifier
@@
 (c) LADM + fixed ETF + Dual regularization
-Figure 5: t-SNE visualisation of test data feature embeddings on the multi-label CIFAR10 experiment with $\beta=0.5$, $\gamma=0.5$.
+Figure 6: t-SNE visualisation of test data feature embeddings on the multi-label CIFAR-10 experiment with $\beta=0.5$, $\gamma=0.5$.
 Each colour represents a class.
@@
 (c) LADM + ETF + Dual regularization
-Figure 6: Pair-wise cosine similarity between test data class-wise average feature prototypes on the multi-label CIFAR10 dataset.
-$\beta=0.5,\gamma=0.5$
-macro-AUC macro-F1 AUC F1
-Centralized (Upperbound) $92.20\pm 0.36$ $61.30\pm 0.94$ $90.04\pm 0.65$ $62.02\pm 0.62$
-FedAvg (McMahan et al.,, 2017) $82.29\pm 0.46$ $39.47\pm 1.59$ $81.48\pm 0.78$ $40.26\pm 1.09$
-FedCurv (Shoham et al.,, 2019) $82.53\pm 0.30$ $39.96\pm 1.28$ $82.10\pm 0.29$ $40.34\pm 1.04$
-FedProx (Li et al.,, 2020) $82.45\pm 0.34$ $39.22\pm 0.74$ $81.77\pm 0.48$ $39.66\pm 0.53$
-SCAFFOLD (Karimireddy et al.,, 2020) $82.51\pm 0.44$ $39.98\pm 1.41$ $82.08\pm 0.53$ $40.26\pm 1.26$
-SphereFed (Dong et al.,, 2022) $\underline{83.63}\pm 1.50$ $42.58\pm 3.20$ $\underline{83.50}\pm 1.87$ $43.18\pm 2.34$
-FedLGT (Liu et al.,, 2024) $83.52\pm 0.68$ $\underline{43.60}\pm 1.68$ $83.36\pm 0.71$ $\underline{44.03}\pm 1.71$
-[…]
-Table 9: Comparisons on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009).
-Non-IID data distributions across clients are simulated using the Dirichlet concentration parameter ($\beta$) and the class-missing ratio ($\gamma$).
-Experiments (additional)
-To complement the results reported in the main manuscript and further validate the reproducibility of our method, we present additional experiments on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009).
-These experiments were repeated three times with different random seeds to evaluate the stability and effectiveness of the proposed approach.
-The mean and standard deviation of the results are reported in Table 9.
-As shown in the table, the proposed method demonstrates stable performance across multiple runs and consistently outperforms competing approaches.
-Dataset
+Figure 7: Pair-wise cosine similarity between test data class-wise average feature prototypes on the multi-label CIFAR-10 dataset.
+To further analyse model behaviour, we visualize the latent feature distributions of the test set using t-SNE under different architectural and training configurations.
+As shown in Figure 6, incorporating LADM for feature disentanglement, along with a predefined ETF classifier to regulate feature distribution across clients, enables the model to focus on semantic content rather than irrelevant factors such as the number of labels present in each sample.
+Furthermore, the addition of regularization terms during training leads to more compact and semantically coherent clusters, while also reducing the similarity among class-wise average prototypes.
+As illustrated in Figure 7, the pairwise cosine similarity between class-wise average features decreases with the inclusion of LADM and the ETF classifier, and is further reduced by the added regularization terms, indicating enhanced inter-class separability and stronger discriminative capability.
+Appendix C Dataset
 In this section, we introduce the datasets used in our experiments.
 To evaluate both the effectiveness and real-world applicability of the proposed method, we conduct experiments on datasets from general computer vision (CV) as well as medical imaging domains.
-Specifically, we use CIFAR-10 (Krizhevsky et al.,, 2009) and PASCAL VOC (Everingham et al.,, 2010) as general CV datasets, and DermaMNIST (Yang et al., 2023a, ) and ChestX-ray14 (Wang et al.,, 2017) as medical datasets.
-We simulate the FL system with 10 clients.
+Specifically, we use CIFAR-10 (krizhevsky2009learning) and PASCAL VOC (everingham2010pascal) as general CV datasets, and DermaMNIST (yang2023medmnist) and ChestX-ray14 (wang2017chestxray) as medical datasets.
+We simulate an FL environment with 10 clients.
 To model non-IID data distributions, the data is partitioned using a Dirichlet distribution, with the concentration parameter $\beta$ controlling the degree of heterogeneity.
-To further simulate missing-class scenarios, we constrain the number of classes available to each client using the class-presenting ratio $\gamma$, which specifies the proportion of total classes present locally.
+To further simulate missing-class scenarios, we constrain the number of classes available to each client using the class presence ratio $\gamma$, which specifies the proportion of total classes present locally.
 Dataset-specific settings are detailed below.
-CIFAR-10 (Krizhevsky et al.,, 2009) is a widely used benchmark in CV.
+CIFAR-10 (krizhevsky2009learning) is a widely used benchmark in CV.
 It comprises 10 classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck.
-To adapt CIFAR-10 for multi-label classification, following Li et al., 2023a , we create composite samples by combining multiple original images into a single image, with their corresponding labels forming a multi-label ground truth.
+To adapt CIFAR-10 for multi-label classification, following li2023neural, we create composite samples by combining multiple original images into a single image, with their corresponding labels forming a multi-label ground truth.
 Specifically, we retain a portion of the original single-label images and augment the dataset by creating an equal number of synthetic samples for each possible pairwise class combination.
-These composites are constructed by randomly selecting two images from different classes and combining them into one.
-We evaluate our method on the multi-label CIFAR-10 dataset under two FL configurations, using $\beta=0.5,\ \gamma=0.5$ and $\beta=0.1,\ \gamma=0.5$.
-Figure 8 illustrates the corresponding class-wise data distribution across clients.
-DermaMNIST (Yang et al., 2023a, ) is a skin lesion classification dataset.
+Each composite is constructed by randomly selecting two images from different classes and merging them into a single sample.
+We evaluate our method on the resulting multi-label CIFAR-10 dataset under two FL configurations: $\beta=0.5,\ \gamma=0.5$ and $\beta=0.1,\ \gamma=0.5$.
+Figure 9 illustrates the corresponding class-wise data distributions across clients.
+Figure 8: Examples from the multi-label DermaMNIST dataset. Each subfigure displays a composite image along with its corresponding set of labels.
+DermaMNIST (yang2023medmnist) is a skin lesion classification dataset.
 It consists of 7 diagnostic categories: actinic keratoses, basal cell carcinoma, benign keratosis-like lesions, dermatofibroma, melanocytic nevi, melanoma, and vascular lesions.
-To adapt DermaMNIST for multi-label classification, we apply a strategy similar to the one used for multi-label CIFAR-10.
+To adapt DermaMNIST for multi-label classification, we apply a strategy similar to that used for the multi-label CIFAR-10 dataset.
 Given the long-tailed nature of the original dataset, we preserve its inherent distribution by retaining all single-label samples.
-Then, we augment the dataset with an equal number of synthetic samples for each possible pairwise label combination.
-Figure 7 shows examples of the resulting composite multi-label inputs.
-The resulting dataset comprises the complete set of original samples along with the newly generated multi-label samples.
+We then augment the dataset by generating an equal number of synthetic samples for each possible pairwise label combination.
+Examples of the resulting composite multi-label samples are shown in Figure 8.
+The final dataset comprises the complete set of original samples together with the newly generated multi-label samples.
 We evaluate our method on this multi-label DermaMNIST dataset under two FL configurations: $\beta=0.5,\gamma=0.71$ and $\beta=0.1,\gamma=0.71$.
 The corresponding class-wise data distributions across clients are illustrated in Figure 10.
-(a) actinic keratoses, vascular lesions
-(b) basal cell carcinoma, dermatofibroma
-(c) bird, vascular lesions
-(d) dermatofibroma, melanoma
-Figure 7: Examples from the multi-label DermaMNIST dataset. Each subfigure displays a composite image along with its corresponding set of labels.
-PASCAL VOC 2012 (VOC) (Everingham et al.,, 2010) is a widely used benchmark dataset in CV.
-It consists of approximately 11,500 images, each annotated with one or more object categories selected from a predefined set of 20 classes.
-These object categories are: aeroplane, bicycle, bird, boat, bottle, bus, car, cat, chair, cow, dining table, dog, horse, motorbike, person, potted plant, sheep, sofa, train, and TV monitor.
-This dataset poses a challenging multi-label classification problem due to the presence of class imbalance, intra-class variability, and frequent inter-class co-occurrence.
-For example, the person category often co-occurs with many other object classes.
-Under a label-skewed FL setting, the non-IID distribution of data and inconsistencies in class co-occurrence patterns across clients further exacerbate the learning difficulty.
-We evaluate our method on VOC under two FL configurations, using $\beta=0.05$, $\gamma=0.5$ and $\beta=0.01$, $\gamma=0.5$.
-Figure 10 illustrates the resulting class-wise data distribution across clients.
-ChestX-ray14 (Wang et al.,, 2017) is a large-scale medical imaging dataset commonly used for automated thoracic disease classification.
-It consists of 112,120 frontal-view chest X-ray images collected from 30,805 unique patients.
-Each image is annotated with one or more disease labels, making the dataset naturally suitable for multi-label classification tasks.
-These labels were extracted from corresponding radiology reports using natural language processing techniques.
+PASCAL VOC (everingham2010pascal) is a widely used benchmark dataset in CV.
+It contains approximately 11,500 images, each annotated with one or more object categories selected from a predefined set of 20 classes.
+These categories include: aeroplane, bicycle, bird, boat, bottle, bus, car, cat, chair, cow, dining table, dog, horse, motorbike, person, potted plant, sheep, sofa, train, and TV monitor.
+This dataset presents a challenging multi-label classification task due to substantial class imbalance, high intra-class variability, and frequent inter-class co-occurrence.
+For instance, the person class commonly appears alongside many other object categories.
+Under a label-skewed FL setting, the non-IID data distribution and inconsistent co-occurrence patterns across clients further exacerbate these challenges.
+We evaluate our method on PASCAL VOC under two FL configurations: $\beta=0.05$, $\gamma=0.5$ and $\beta=0.01$, $\gamma=0.5$.
+Figure 11 illustrates the resulting class-wise data distributions across clients.
+ChestX-ray14 (wang2017chestxray) is a large-scale medical imaging dataset widely used for automated thoracic disease classification.
+It contains 112,120 frontal-view chest X-ray images collected from 30,805 unique patients.
+Each image is annotated with one or more disease labels, making the dataset naturally suited to multi-label classification.
+These labels are extracted from the corresponding radiology reports using natural language processing techniques.
 The dataset includes 14 disease categories: Atelectasis, Cardiomegaly, Consolidation, Edema, Effusion, Emphysema, Fibrosis, Hernia, Infiltration, Mass, Nodule, Pleural Thickening, Pneumonia, and Pneumothorax.
 In addition, a No Finding label is used to indicate negative samples in which none of the 14 diseases are present.
-Due to its scale, real-world variability, and label noise, ChestX-ray14 provides a challenging benchmark for developing and evaluating deep learning models in medical image analysis, particularly in multi-label, imbalanced scenarios.
-We conduct experiments under two label-skewed FL settings on this dataset, using configurations of $\beta=0.5$, $\gamma=0.5$ and $\beta=0.1$, $\gamma=0.5$.
-Since a significant portion of the dataset is “No Finding” samples, we distribute these evenly across all clients in both settings.
+Due to its scale, real-world variability, and inherent label noise, ChestX-ray14 provides a challenging benchmark for developing and evaluating deep learning models in medical image analysis, particularly in multi-label and imbalanced scenarios.
+We conduct experiments on this dataset under two label-skewed FL configurations: $\beta=0.5$, $\gamma=0.5$ and $\beta=0.1$, $\gamma=0.5$.
+Since a significant portion of the dataset (57% of the training data) is No Finding samples, we distribute these samples evenly across all clients in both settings.
 This setup mimics realistic clinical scenarios where healthy cases are common, while disease cases are relatively rare and unequally distributed across institutions.
-Figure 11 shows the class-wise data distribution across clients.
+Figure 12 illustrates the resulting class-wise data distributions across clients.
 (a) Overall training data distribution.
@@
 (f) Per-client training data bubble chart under $\beta=0.1$ and $\gamma=0.5$.
-Figure 8: Distribution of data across local clients used in the experiments on the CIFAR-10 dataset (Krizhevsky et al.,, 2009).
+Figure 9: Distribution of data across local clients in the CIFAR-10 (krizhevsky2009learning) experiments.
+The class presence ratio ($\gamma$) is set to 0.5 ($\leq$ 5 of 10 classes per client).
+Non-IID client distributions are simulated using the Dirichlet factor ($\beta$).
 (a) Overall training data distribution.
@@
 (f) Per-client training data bubble chart under $\beta=0.1$ and $\gamma=0.71$.
-Figure 9: Distribution of data across local clients used in the experiments on the DermaMNIST dataset (Yang et al., 2023a, ).
+Figure 10: Distribution of data across local clients in the DermaMNIST (yang2023medmnist) experiments.
+The class presence ratio ($\gamma$) is set to 0.71 ($\leq$ 5 of 7 classes per client).
+Non-IID client distributions are simulated using the Dirichlet factor ($\beta$).
 (a) Overall training data distribution.
@@
 (f) Per-client training data bubble chart under $\beta=0.01$ and $\gamma=0.5$.
-Figure 10: Distribution of data across local clients used in the experiments on the VOC dataset (Everingham et al.,, 2010).
+Figure 11: Distribution of data across local clients in the PASCAL VOC (everingham2010pascal) experiments.
+The class presence ratio ($\gamma$) is set to 0.5 ($\leq$ 10 of 20 classes per client).
+Non-IID client distributions are simulated using the Dirichlet factor ($\beta$).
 (a) Overall training data distribution.
@@
 (f) Per-client training data bubble chart under $\beta=0.1$ and $\gamma=0.5$.
-Figure 11: Distribution of data across local clients used in the experiments on the ChestX-ray14 dataset (Wang et al.,, 2017).
+Figure 12: Distribution of data across local clients in the ChestX-ray14 (wang2017chestxray) experiments.
+The ChestX-ray14 dataset contains 14 thoracic disease categories and an additional “No Finding” label.
+Since a large portion of the dataset (57% of the training data) is “No Finding” samples (i.e., negative cases with all-zero labels), we distribute these samples evenly across all clients to reflect a realistic clinical scenario in which healthy cases are prevalent.
+The class presence ratio ($\gamma$) is set to 0.5 ($\leq$ 7 of 14 disease classes per client).
+Non-IID client distributions are simulated using the Dirichlet factor ($\beta$).